### v1

# Modular Compilation for Quantum Chiplet Architectures

Abstract.
As quantum computing technology continues to mature, industry is adopting modular quantum architectures to keep quantum scaling on the projected path and meet performance targets. However, the complexity of chiplet-based quantum devices, coupled with their growing size, presents an imminent scalability challenge for quantum compilation. Contemporary compilation methods are not well-suited to chiplet architectures — in particular, existing qubit allocation methods are often unable to contend with inter-chiplet links, which don’t necessary support a universal basis gate set. Furthermore, existing methods of logical-to-physical qubit placement, swap insertion (routing), unitary synthesis, and/or optimization are typically not designed for qubit links of wildly varying levels of duration or fidelity. In this work, we propose SEQC, a complete and parallelized compilation pipeline optimized for chiplet-based quantum computers, including several novel methods for qubit placement, qubit routing, and circuit optimization. SEQC attains up to a 36% increase in circuit fidelity, accompanied by execution time improvements of up to 1.92×1.92$\times$1.92 ×. Additionally, owning to its ability to parallelize compilation, SEQC achieves consistent solve time improvements of 24×2-4$\times$2 - 4 × over a chiplet-aware Qiskit baseline.
Quantum computing, quantum compilation, modular architectures.
copyright: nonejournalyear: 2024doi: XXXXXXX.XXXXXXX
1. Introduction
Classical computer systems, through the decades of their existence, have become increasingly distributed. Physical, technological, and economic constraints have prevented single “monolithic” systems from scaling to the point that they could meet the demand for high performance and yet remain practical. Today, distributed architectures are prevalent, and their latest renditions, from cloud computing to chiplet-based digital processors, are ubiquitous.
We postulate that quantum computing is on a similar path. While improvements in superconducting quantum hardware have led to the debut of processors with $1000+$ qubits (ibm_condor), practical implementations of quantum computation will require millions of physical qubits (paler2019really). The number of qubits required by a quantum algorithm, the depth of the algorithm (i.e., execution time, or number of gates on the critical path), the auxiliary and syndrome qubits required to reach sufficiently low error rates, and the size of matter qubits when all control hardware is included, all suggest that many more qubits are required than are likely to fit in a single die (vanMeter_2010_DQC). The overhead for error correction, cooling, dilution, control systems, I/O lines, challenges associated with verification and adequate chip thermalization, and realistic resources such as lossy waveguides, limited qubit fabrication yields as qubit capacity grows (smith_2022_chiplets), and finite chip sizes, make the prospect of a single “monolithic” quantum processor very expensive or entirely unrealistic by current standards. These constraints force large-scale quantum systems to adopt a physically distributed architecture.
We are already observing signs of this shift. Recent developments in quantum chip linking, including flip-chip architectures (gold2021entanglement) and low-loss coaxial cables (niu2023low), suggest that modular designs are the most viable for scaling quantum computers (smith_2022_chiplets). Many contemporary or upcoming leading quantum systems adopt chiplet-based modular quantum processors, for example using high-bandwidth quantum links between nearest-neighbor quantum chiplets (e.g., carrier-chip couplers in Rigetti Aspen-M (Rigetti_AspenM; gold2021entanglement), or m-couplers in IBM Crossbill (gambetta_2024_IBMcouplers; Mandelbaum_2024_IBMCouplers)), or lower-fidelity, lower-bandwidth, but longer-distance flexible coupling of discrete chips (e.g., l-couplers in IBM Flamingo (gambetta_2024_IBMcouplers; Mandelbaum_2024_IBMCouplers)), or a combination of the above (e.g., c-, m-, and l-couplers in IBM Starling (Mandelbaum_2024_IBMCouplers)), or multi-chip connectivity through tunable couplers and routing chips (Field_2024_RigettiModular). Most major companies have set their sights at modular designs to meet quantum scaling targets in practical ways, and modular quantum processors dominate their latest roadmaps (IBM_2023_Roadmap; IonQ_2024_Roadmap; Rigetti_2024_Roadmap).
Unfortunately, current compilation infrastructure is not capable of reasoning well about this quantum interconnect heterogeneity. The mapping of a quantum program’s logical quantum gates into the native gates supported by the underlying quantum processor, and the mapping of logical qubits to the physical qubits of the quantum hardware, largely determine the number and type of native quantum gates required for the computation, the circuit depth, and its execution time—long programs may not complete successfully as qubits decohere. Different mappings can result in vastly different compiled quantum circuits, with diverse characteristics along all these axes. These details of efficiency can make or break an algorithm in today’s noisy intermediate-scale quantum systems (NISQ) era, so quantum compilers aggressively optimize for all of them. To make matters more challenging, each qubit, coupler, and gate have diverse error profiles that are highly variable both spatially and temporally (Tannu_2019_ErrorVariability; Murali_2019_NoiseMapping; dasgupta2021stability). Unlike classical compilation that is done only once for a given architecture, quantum programs must be recompiled every time before execution, as the ideal physical qubits to execute on change between runs. So, quantum compilers today are faced with the daunting task of optimizing quantum programs across multiple dimensions with often conflicting demands, in a continuously changing environment. Coupled with the hardness of synthesizing unitaries (which is exponential in the number of qubits) and the need for quick and frequent compilation, modern quantum compilers have no option other than to rely on heuristics to perform the task. These heuristics result in compilation complexity of $O(n^{2})$ for $n$ total qubits in a quantum processor.
Hardware modularity adds significant complexity to this already hard task, as inter-chiplet links are typically inferior compared to intra-chiplet ones (laracuente2022modeling; smith_2022_chiplets), connectivity across chiplets is often limited (gambetta_2024_IBMcouplers; Mandelbaum_2024_IBMCouplers), and not all basis gates are necessarily supported across chiplets (niu2023low). In fact, popular quantum software stacks today (e.g., Qiskit (qiskit2024)) are not even cognizant of the existence of hardware modularity. Hardware modularity, though, also presents an opportunity. Inspired by classical compilation, in this paper we leverage hardware modularity to achieve compilation modularity.
Compilation in classical systems is typically performed independently and in parallel for each source file, producing one object file per source. The individual object files are then linked together to construct the executable. We propose a compilation framework for modular quantum processors that works in a similar fashion: it stratifies, i.e., splits, the source quantum circuit into subcircuits small enough to fit in each chiplet, and maps subcircuits to chiplets, and then in parallel elaborates each subcircuit and compiles it for its target chiplet. This Stratify-Elaborate Quantum Compiler (SEQC) stratifies a source program only once for a given chiplet architecture, and performs only the elaboration step recurrently before each execution. In essence, SEQC replaces the recurrent $O(n^{2})$ compilation step for an $n$-qubit quantum processor, with several parallel $O(k^{2})$ elaboration steps for $k$-qubit chiplets. As the qubit capacity $n$ of quantum processors grows exponentially, today’s $O(n^{2})$ compilation latency rises even faster. We expect, however, that the number of qubits $k$ per chiplet will remain relatively stable or grow much slower, as it seems to be the case for the foreseeable future (IBM_2023_Roadmap), and thus the SEQC recurrent compilation latency is expected to remain relatively stable. The stratification step is $O(n^{2})$, so the end-to-end complexity remains the same, but stratification is performed only once; the recurrent compilation in SEQC is only $O(k^{2})$, and barely growing with new processor designs.
Additionally, as SEQC is cognizant of hardware modularity, it can stratify the source quantum circuit into subcircuits and map them to chiplets to minimize inter-chiplet communication. As we show in this paper, SEQC produces circuits with shorter execution times and significantly fewer inter-chiplet gates compared to today’s stock compilers, leading to much higher fidelity execution. More importantly, as the number of qubits in a processor grows, SEQC achieves even higher performance in these figures of merit.
In summary, the contributions of this paper are as follows:
  • We make stock compilers aware of hardware modularity, thereby allowing them to correctly compile circuits for modular architectures with limited cross-chiplet gate support.
  • We design and implement SEQC, a Stratify-Elaborate Quantum Compiler for modular architectures. SEQC performs compilation in two stages, with the first stage (stratification, or chiplet splitting) performed only once for a given architecture, and the second stage (elaboration, or chiplet compilation) performed in parallel for each chiplet. Only this second stage needs to be performed before each execution, and thus SEQC’s compilation time is largely unaffected by the growth of qubit counts in future quantum processors.
  • We design and implement in SEQC several novel methods for qubit placement, qubit routing, and circuit optimization.
  • We evaluate SEQC and show it compiles circuits with up to 36% higher circuit fidelity and up to 1.92×1.92$\times$1.92 × lower execution time, while consistently achieving 24×2-4$\times$2 - 4 × faster compilation time compared to a chiplet-aware Qiskit baseline.
2. Background and Motivation
To execute a program on a target device, the logical program representation must be transformed such that it conforms to the idiosyncratic physical constraints for that particular device. This process of logical-to-physical translation represents a critical part of compilation and is present even in exotic computing technologies, e.g., quantum. In quantum computing, programs are (usually) represented as quantum circuits, and devices are subject to domain-specific physical constraints, such as decoherence time (mike_and_ike), qubit topology (the connectivity between physical qubits) (qubit_allocation; qubit_routing), and the set of basis gates (the family of operations physically implemented on the device). Thus, quantum circuit compilation is necessary to actually run quantum programs on real-world devices.
Compilation is often divided into distinct stages such as qubit allocation (layout and routing), basis translation, and optimization. The relations between each stage and the physical constraints of both monolithic and chiplet architectures are described below.
2.1. Qubit Allocation
Qubit allocation addresses the topology constraints of quantum devices, where physical qubits on a device have limited connectivity to other qubits on the device. It is critical that after this stage, which often requires the modification or addition of circuit gates, the resulting optimized and technology-dependent circuit is functionally equivalent to the original, technology-independent algorithm (smith2019quantum). The qubit allocation problem has been previously shown to be NP-complete (qubit_allocation); thus, algorithms often adopt various strategies to make the problem more tractable. One common method is to further divide qubit allocation into two substages: layout and routing.
Logical-to-physical qubit layout, also known as qubit placement, involves deriving an injective map from the virtual qubits of the quantum circuit to the physical qubits of the quantum device (qubit_routing; sabre_swap). In other words, the qubit layout stage decides the initial position of logical qubits on the device. In contrast, qubit routing operates from an initial qubit layout and inserts SWAP operations to align a quantum circuit’s gates to a given qubit topology. Consider, for example, the state-of-the-art in qubit allocation: the SABRE heuristic search algorithm (sabre_swap). In SABRE, the distinction of the layout and routing stages motivates an iterative method for solving qubit allocation. Namely, an initial layout is used to perform qubit routing, which is used to generate a new initial layout.
Because qubit routing remains NP-hard (qubit_routing), another method of simplifying the problem is to assume qubit links are roughly equivalent, particularly with respect to average fidelity and two-qubit gate duration. This assumption can manifest implicitly when a given device topology is represented as an unweighted, undirected graph. For monolithic architectures, one could debate the marginal tradeoff between performance and solution quality; however, for modular architectures, the cross-chiplet connections are far worse compared to intra-chiplet connections (smith_2022_chiplets). Thus, we could expect the solutions generated from the simplified model to be worse than a more physically accurate model, for example, in circumstances where the shortest path in terms of the quantity or depth of inserted SWAPs diverges from the shortest path in terms of fidelity, gate duration, or some combination thereof.
2.2. Basis Translation
Basis translation converts the gates of an input quantum circuit into gates that are physically supported on hardware. Monolithic quantum architectures typically support a so-called universal basis gate set, where any unitary quantum operation can be constructed using solely the gates in the set (mike_and_ike). Moreover, it has been shown that basis translation can be performed efficiently given the target gate set is universal, and the quantum circuit contains a fixed, finite number of qubits (Solovay–Kitaev). Unfortunately, modular quantum architectures depend upon specialized links to connect otherwise-independent chiplets, and these links often don’t support universal operations (niu2023low). Thus, when algorithms like SABRE depend on the expectation of universality among qubit links, they may generate incompatible outputs when targeting modular architectures.
2.3. Optimization
Quantum circuit optimization seeks to prune extraneous operations by finding opportunities to combine, eliminate, and/or parallelize quantum gates (optimization1). Through commutation and cancellation identities, it is possible to search for optimization opportunities by iteratively transforming the quantum circuit without affecting the underlying operation (optimization1). However, in general, optimization is highly time- and resource-intensive, with the problem scaling exponentially with respect to number of qubits and/or circuit depth (qfactor). As a result, partitioning, the division of a quantum circuit into blocks of quantum gates, is a common tactic among optimization techniques (qfactor; qgo; bqskit). One major consideration in this process is demarcation, i.e., where and when to draw the line between blocks. Namely, one forfeits all optimization opportunities along boundaries of demarcation, so they must be carefully chosen to minimize negative impact on the output solution. From this perspective, modular architectures are naturally synergistic with partitioning, since cross-chiplet connections present a physically motivated boundary between operations.
3. A Strawman Chiplet Compiler
(a) A physical inter-chiplet link exists between halo qubits $A_{0}$ and $B_{0}$. A physical intra-chiplet link exists between $B_{0}$ and $B_{1}$.
(b) A logical CNOT gate between chiplets $A$ and $B$ which is not implementable on hardware (indicated in red).
(c) To prepare for an inter-chiplet SWAP, the logical qubit on $B_{0}$ must be moved to a nearest neighbor, $B_{1}$ in this case, with a SWAP operation.
(d) The inter-chiplet SWAP between chiplets $A$ and $B$ completes transformation of the CNOT gate to an intra-chiplet gate.
(e) Once the CNOT gate has been successfully placed, the inserted SWAPs must be undone to not affect future gates in the circuit.
Figure 1. Peephole correction for incompatible (non-SWAP) gates placed on inter-chiplet halo edges.
It was challenging to establish an appropriate baseline for state-of-the-art quantum circuit compilation on chiplet-based quantum devices. Not only are modular quantum architectures relatively nascent, but compilation schemes for monolithic devices are largely incompatible with chiplet devices, as was discussed in greater detail in Section 2.
Thus, we took it upon ourselves to design a baseline compilation scheme for chiplet architectures, against which we compare our more-optimized design presented in Section 4. This “strawman” compiler augments existing compiler designs with the minimal modifications necessary to generate valid circuit outputs for chiplet architectures. Specifically, we insert a “peephole” compiler pass immediately after qubit routing, which corrects for the placement of two-qubit gates on inter-chiplet links as shown in Figure 1. This algorithm is inefficient, since it inserts a total of 4 SWAP gates — 2 inter-chiplet and 2 intra-chiplet — for every inter-chiplet gate. Unfortunately, this degree of overhead is inevitable when the stock, unmodified compiler is ignorant of chiplet-specific device constraints.
4. The Design of SEQC
We incorporate the hierarchical elements of chiplet architectures into our compiler design by adopting a two-stage procedure. The first, stratification stage, represents a one-time overhead to elevate a quantum circuit to the chiplet level. The second, elaboration stage, represents a recurring cost to complete the compilation of a quantum circuit given the most up-to-date backend properties.
4.1. Stratification Stage
(a) A $3$-chiplet topology with $2$ qubits per chiplet ($6$ qubits in total).
(b) Mapping of a $6$-qubit quantum circuit to $3$ subcircuits. Gates represent arbitrary cross-subcircuit operations.
(c) Mapping of subcircuits to physical chiplets. Since chiplets $A$ and $C$ are not physically connected, certain gates (shown in red) necessitate inter-chiplet routing.
(d) Inter-chiplet routing is performed to address cross-chiplet gates that either don’t conform to the physical inter-chiplet topology or the basis gate set (aren’t SWAP gates).
Figure 2. Showcase of Inter-Chiplet Compilation in the Stratification Stage
In this stage, we map a quantum circuit to the inter-chiplet device topology, e.g., the structure in Figure 2(a), by stratifying it into smaller, interconnected subcircuits. This can be accomplished with the following intermediate steps, as shown in Figures 2(b)2(d). First, qubit-to-subcircuit mapping designates which logical qubits belong to which logical subcircuits. Then, as part of chiplet allocation, we define a mapping of logical subcircuits to physical chiplets on the target device and route inter-chiplet gates to conform with constraints of physical-chiplet connectivity.
4.1.1. Qubit-to-Subcircuit Mapping
To perform qubit-to-subcircuit mapping (Figure 2(b)), we leverage a form of simulated annealing. Simulated annealing is a form of stochasic optimization, notably used in VLSI algorithms for place-and-route (annealing1989). We use simulated annealing to find a qubit assignment that minimizes the number of connections between subcircuits. As our chiplet allocation algorithm acts conservatively—performing the minimal number of changes necessary to make a circuit valid—it will produce better results if given a circuit with minimal connectivity between subcircuits. For our cost function, we count the number of inter-subcircuit gates. To permute the solution, we swap two random qubits in different subcircuits. As different simulated annealing runs are independent, running the algorithm using different seeds is embarrassingly parallel. We exploit this parallelism by running many trials at once, each in a different process with a different seed, then choosing the best result across all trials and seeds.
4.1.2. Chiplet Allocation
To perform chiplet-level allocation, shown in Figures 2(c) and 2(d), we extend and modify the SABRE algorithm for initial mapping (sabre_swap) with application-specific layout and routing passes. For layout, we initialize each iterative allocation cycle with a random placement of logical subcircuits into physical chiplets. For routing, we developed a new heuristic for deciding between candidate SWAPs.
By default, SABRE attempts to place two-qubit gates on a valid edge in the graph of a device’s hardware topology, i.e, the physical distance between the qubits of a gate should be $1$. For inter-chiplet routing, we do not wish to factor intra-qubit topology into our distance calculation. Furthermore, we hope to move all gates to be within the same chiplet, i.e., the distance between qubits should be 00 instead of $1$. Thus, when selecting between potential SWAPs, we categorize and prioritize SWAPs in a three-tier approach:
  1. (1)
    Symbiotic SWAPs: A SWAP between two multi-chiplet gates is considered symbiotic if it reduces the distance of both gates, i.e., each gate benefits from the SWAP. In theory, we should expect symbiotic SWAPs to be the most efficient, and thus highest-priority, SWAP possible.
  2. (2)
    Commensalistic SWAPs: When a SWAP can benefit one multi-chiplet gate without harming any other, we consider it to be commensalistic. Such behavior occurs most commonly on idle qubits in the quantum circuit, although it can occur between two multi-chiplet gates.
  3. (3)
    Parasitic SWAPs: The remaining category of SWAPs to consider benefit one gate while harming another. Since this variety of SWAP is actively counterproductive towards a gate, it should be, and likely will be, avoided in most circumstances. However, given a particularly onerous circuit, it may be necessary in such rare instances. For example, clear one gate may free physical qubits to help route others.
With these changes, our algorithm is able to correctly and efficiently place and route circuits for any valid device topology.
4.2. Elaboration Stage
(a) A $3$-qubit chiplet with a line topology and connected to a neighboring chiplet by $q_{2}$.
(b) A $3$-qubit subcircuit with $2$ inter-chiplet SWAPs.
[…]
(d) A greedy policy takes the shortest path for routing inter-chiplet gates to their nearest halo qubit.
(e) Qubit routing for the remaining intra-chiplet gates is performed as normal, except with restrictions to ensure inter-chiplet SWAPs are kept static.
(f) The remaining compilation states (basis translation and optimization) are not affected by the additional constraints imposed by inter-chiplet gates.
Figure 3. Showcase of Intra-Chiplet Compilation in the Elaboration Stage.
Following the stratification stage, each chiplet can be compiled almost completely independently from one another. More precisely, the compilation process within each chiplet is largely identical to that of a monolithic quantum architecture. The only exception is the presence of inter-chiplet SWAPs, which must be lowered from the chiplet level to the qubit level. Once this assignment is made, these SWAPs must remain fixed and immutable to subsequent compilation stages. Despite this overhead, restricting the task of compilation to a narrower domain of qubits facilitates both parallelization (as each subcircuit can be processed in parallel) and problem size reduction (as there are less qubits to consider) of the most computationally intensive compilation stages (routing and optimization). To accomplish these tasks, we divide the elaboration stage into two substages: qubit allocation and parallel basis translation and optimization. Figure 3 illustrates this with an example of elaboration for a single chiplet in Figure 3(a) and its corresponding subcircuit in Figure 3(b).
4.2.1. Qubit Allocation
Like other qubit allocation schemes, we incorporate layout (see Figure 3(c)) and routing (see Figure 3(e)) stages. The presence of inter-chiplet gates necessitates several key changes to the SABRE (sabre_swap) qubit allocation algorithm to achieve correctness and/or maximum parallelism. First, the placement of inter-chiplet gates is a non-parallelizable task (due to cross-chiplet interactions) that strongly influences the layout and routing passes, which could otherwise be fully parallelized across chiplets. To account for this, we introduce a serial stage (shown in Figure 3(d)) between the parallelized layout and routing passes that greedily places inter-chiplet gates on the nearest valid edge between halo qubits. A greedy policy was chosen because inter-chiplet gates have more restrictions on valid placements compared to intra-chiplet gates, while also experiencing substantially higher error rates (smith_2022_chiplets). Once the positions of the inter-chiplet gates are settled, they must remain fixed in place, imposing an additional constraint on qubit routing. To promote this behavior in SABRE, we extended its cost heuristic to weigh the device topology by fidelity. This causes higher error links to assume higher costs, discouraging SABRE from moving the high-error inter-chiplet gates. As a result, SEQC achieves better solutions in the common case. However, there is nothing to prevent SABRE from incorrectly modifying any inter-chiplet SWAPs. Thus, we need to impose hard restrictions to ensure correctness. We establish two disjoint gates sets. One gate set operates strictly within a chiplet. This is the traditional gate set for a monolithic processor. The other gate set is the only one permitted to operate across chiplets, and specifically only implements inter-chiplet SWAPs, i.e., SWAP gates between two halo qubits. By restricting each gate set to its own disjoint part of the overall device topology, we ensure correctness across all cases.
4.2.2. Parallel Translation and Optimization
The remaining compilation stages of basis translation and circuit optimization can be easily parallelized at the chiplet granularity, as shown in Figure 3(f). Basis translation, in particular, is embarrassingly parallel, as every gate can be translated independently from every other gate. Meanwhile, as discussed in Section 2, optimization algorithms already employ partitioning and parallelization to create tractable subproblems from intractably large circuits, facilitating a smaller problem space, lower memory usage, and faster performance. In modular quantum architectures, we physically motivate partitions through the natural chiplet boundaries, enabling us to reap the same resource and performance benefits of partitioning without compromising on solution quality.
5. Methodology
To experimentally validate our compiler for modular quantum architectures, we ran a benchmark suite on a wide spectrum of circuit sizes for both SEQC and the strawman baseline. Both SEQC and the baseline compiler were implemented in the Qiskit SDK (qiskit2024) v1.2.41.2.41.2.41.2.4 and Python $3.12.6$. Elaboration time results were taken on a machine with a $128$-core Ampere Altra Max ARM64 processor and $256$ GB of DDR4 RAM. When given a parallelizable workload, SEQC attempts to allocate its tasks evenly among the available physical cores, rounded to the neared whole number. Finally, the simulated annealing technique used in the qubit-to-subcircuit stage uses a starting temperature $T_{0}=200$K, a rate of change of $0.005T$ each round, $\frac{T}{50}$ permutations per round, and a total of $5$ trials per core.
5.1. Metrics
We compare SEQC to the strawman baseline over a wide range of common and/or relevant metrics, discussed below. We aggregate results in a manner similar to the SPEC family of benchmarks (spec2017), by taking the geometric mean of the relative performance ratio against a baseline technique across a benchmark suite of quantum circuits. The resultant score is then derived for each of the evaluated metrics on multiple backends of varying numbers of chiplets.
Fidelity. The top priority for quantum circuit execution is to maximize fidelity. Without a guarantee of correctness, the outputs of the quantum circuit are useless, and any speedup or quantum advantage is rendered moot. Fidelity is impacted by the decoherence constraints of individual qubits, as well as the errors on a gate-by-gate basis (mike_and_ike). Due to simulation and hardware limitations, we utilize the estimated success probability (ESP) metric (esp) to approximate fidelity. Specifically, we track the per-qubit ESP and report an aggregate ESP for each benchmark from the geometric mean of the accumulated measurements.
Non-recurring stratification time. This represents the time taken in the stratification stage, which will be present for our compiler and absent for the baseline. The stratification time represents a one-time overhead to prepare a quantum circuit for execution on a modular architecture. This result can be reused for future executions on that device or any device with an equivalent or more expansive topology.
Recurring elaboration time. The error characteristics of a quantum device can change on a hour-by-hour basis, or even a run-to-run basis (Tannu_2019_ErrorVariability; Murali_2019_NoiseMapping; dasgupta2021stability). Thus, to generate the highest-quality circuits with the most up-to-date characteristics possible, a given quantum circuit should be recompiled for every execution. Since the qubit and chiplet topologies remain static per-device, it is possible under the SEQC framework to reuse the results from the stratification stage, and only incur overhead from repeating the elaboration stage.
Solve time. This metric is relevant only to the baseline strawman chiplet compiler, and represents the time it takes to compile a quantum circuit. The baseline compiler solve time is a recurring cost, i.e., this compilation step has to be performed every time a circuit needs to execute on a quantum machine.
Estimated execution time. The per-shot execution time of a quantum circuit is determined by the critical path of gate operations weighted by gate duration. In addition to performance concerns, execution time indirectly affects fidelity on NISQ-era devices due to decoherence. In other words, it is imperative that quantum circuits execute within the decoherence time to ensure high-fidelity outcomes.
Number of inter-chiplet gates. The inter-chiplet gates of a chiplet architecture contribute the most to error and execution time relative to other gates (at least by a factor of 4×4$\times$4 × (smith_2022_chiplets)). As a result, the number of inter-chiplet gates should represent a strong proxy metric for the quality of a compiler solution. Accordingly, minimizing the number of inter-chiplet gates should result in higher fidelity, faster-executing quantum circuits.
Circuit depth. Defined as the unweighted critical path of a quantum circuit (mike_and_ike), the circuit depth is often employed as a backend-agnostic metric for execution time that can be measured statically. Notably, it is used in SABRE (sabre_swap) as part of its cost heuristic.
Gate count. Similar to circuit depth, the total number of gates in a circuit can also be used as a static, backend-agnostic metric for certain properties of the aforementioned circuit. For instance, it may provide an indication of the amount of accumulated error, or, like circuit depth, the execution time. SABRE (sabre_swap) uses gate count to decide between randomly seeded allocation trials.
5.2. Benchmarks
For our benchmark suite, we leverage a subset of circuits in the Supermarq (supermarq) suite of quantum circuits. The circuits we select are chosen to be representative of practical, real-world applications for quantum computing, while also being feasible to compile for a large number of qubits. In particular, our benchmark suite comprises the following quantum circuits.
[…]
VQE. The variational quantum eigensolver (VQE) (vqe) is another hybrid quantum-classical variational algorithm that is often used in chemistry applications. In our experiments, we choose a $2$-layer ansatz whose parameters are once again randomly generated.
Hamiltonian Simulation. Quantum devices are naturally well-suited to simulating Hamiltonians. The Supermarq benchmark selects the 1D Transverse Field Ising Model (TFIM) system to model, which is relevant for phase transitions in magnetic materials.
5.3. Simulated Quantum Device Specifications
We perform our experiments on mock backends that are generated to conform to the chiplet architecture proposed in Smith et al. (smith_2022_chiplets). The hardware specifications of the qubits and intra-chiplet connections are sourced from Acharya et al. (backend_specs). Table 1 details the full specifications. To generate the expected fidelity of the inter-chiplet SWAPs, we perform a simulated random benchmarking trial and calculate the error rate of a SWAP gate with these backend specifications. Following Smith et al. (smith_2022_chiplets), we model inter-chiplet SWAPs on this backend with 4×4$\times$4 × the intra-chiplet error rate. Similarly, we approximate the duration of the inter-chiplet SWAP gate as 4×4$\times$4 × the duration of an intra-chiplet SWAP.
T1 (backend_specs) $20\times 10^{-6}$ seconds
T2 (backend_specs) $30\times 10^{-6}$ seconds
Frequency (backend_specs) $6\times 10^{9}$ Hz
(a) Qubit Properties.
Instruction Duration (ns) Error
(backend_specs) $25$ $0.109\%$
SX (backend_specs) $25$ $0.109\%$
$\text{R}_{\text{z}}(\phi)$ (backend_specs) 00 $0.00\%$
CZ (intra-chiplet) (backend_specs) $34$ $0.605\%$
SWAP (inter-chiplet) $702.4$ $10.23\%$
Reset (backend_specs) $500$ $0.186\%$
Measure (backend_specs) $500$ $0.196\%$
(b) Instruction Properties.
Table 1. Specifications of Simulated Chiplet Backend.
The topology of the simulated device is composed of a grid lattice of symmetric chiplet modules, where each chiplet features a heavy-hexagon qubit topology (heavy_hex), as shown in Figure 4. The intra-chiplet topology is constructed such that the heavy-hex structure is preserved for the global device topology (ignoring chiplet division). The tessellation requirements of each chiplet restrict valid chiplet sizes to multiples of $10$; however, for our experiments, we limit the size of each chiplet to the minimum of $10$ qubits due to hardware limitations.
Note that for each valid chiplet count, there may exist multiple valid chiplet topologies. In this work, we generate our backends to use the “most-square” topology for a given chiplet count. For example, a $12$-chiplet machine would be generated with a grid of $3\times 4$ chiplets.
Figure 4. Construction of a grid inter-chiplet lattice from a heavy-hexagon qubit lattice, where the heavy-hex lattice is maintained for each chiplet.
6. Experimental Results
Figures 57 present measurements of the recurring compilation time (i.e., solve time) of the baseline strawman chiplet compiler, and SEQC’s two compilation stages: the non-recurring stratification stage and the recurring elaboration stage.
Figure 5. SEQC stratification time normalized to the 20-qubit (i.e., 2-chiplet) trial.
The average stratification time of SEQC across the benchmark suite obeys a quadratic trajectory ($R^{2}=0.9971$) with respect to the circuit size and number of chiplets, as shown in Figure 5. We note that the compilation time of the baseline, presented in Figure 6, also scales quadratically ($R^{2}=0.9964$). Although stratification time represents a one-time cost, the quadratic scaling of stratification points toward the practicality of SEQC. The stratification time depends heavily on the complexity of the input circuit. Simple circuits, such as GHZ or HamiltonianSimulation, exhibit a very gradual increase in stratification time as circuit size increases, while others, such as BitCode and PhaseCode, are much steeper. This observation holds true for the solve time of the baseline compiler as well. Collectively, these results indicate that the computational complexity of the baseline compiler and the stratification stage of SEQC are affected similarly by the same quantum circuit characteristics.
Figure 6. Baseline compilation time normalized to the 20-qubit (i.e., 2-chiplet) trial.
Figure 7. Relative elaboration time of SEQC normalized to the strawman baseline for devices with 10-qubit chiplets (lower is better).
From Figure 7, we observe an improved SEQC elaboration time compared to the baseline solve time, leading to a roughly 24×2-4$\times$2 - 4 × speedup, up until $800$-qubit (i.e., $80$-chiplet) trials. The observed speedup reflects the expected performance improvement that our stratify-elaborate technique should garner from parallelization and problem-space reduction. For $81$-chiplet experiments and above, our results begin to reflect a far more marginal speedup compared to baseline, reflecting the limitations of our evaluation hardware. Namely, as the number of chiplets begins to approach the number of physical cores and hardware threads on our hardware, SEQC experiences difficulties in evenly distributing its tasks. For example, under-allocating tasks causes poor core utilization, while over-allocation leads to process thrashing, either way resulting in poorer elaboration time. However, in all cases, SEQC outperforms the baseline on average.
On a circuit-by-circuit basis, we notice higher relative performance with parallel workloads (like BitCode and PhaseCode) compared to serial workloads (GHZ and HamiltonianSimulation). Notably, above the $81$-chiplet threshold where the classical machine we use for compilation no longer suffices, elaboration time exceeds the baseline solve time for the serial workloads. This is because serial circuits are fundamentally trivial to solve, and thus benefit less from the parallelism offered by SEQC. We verified experimentally that in these cases, SEQC spends almost all of its time on thread startup and teardown, rather than useful work. Moreover, the raw compilation time is trivially small in these cases, owning to the simplicity of the circuits, thus relative performance differences have little practical significance.
Figure 8. Relative estimated fidelity (ESP) of SEQC normalized to the strawman baseline for devices with 10-qubit chiplets (higher is better).
Figure 8 presents the relative estimated fidelity of SEQC compared to the baseline with respect to the number of qubits per circuit. Notably, as the number of chiplets / the circuit size increases, SEQC improves its relative fidelity against the baseline. We begin to see higher fidelity circuits in the geometric mean once the circuit size exceeds $36$ chiplets, up to a maximum of $36\%$ higher fidelity. All benchmarks reflect higher relative performance for the $50$-chiplet trials and above.
Figure 9. Estimated per-shot execution time of circuits produced by SEQC normalized to circuits produced by the strawman baseline for devices with 10-qubit chiplets (lower is better).
Figure 9 presents the estimated per-shot execution time of the circuits produces by SEQC compared to the circuits produced by the baseline compiler, as a function of the number of qubits per circuit. As explained above, we approximate the duration of an inter-chiplet SWAP as 4×4$\times$4 × the duration of an intra-chiplet SWAP (smith_2022_chiplets). Under this model, we observe execution time benefits up to a 1.92×1.92$\times$1.92 × improvement over baseline at $750$ qubits ($75$ chiplets). Not only are improvements in execution time desirable for quantum circuit throughput, they could also result in lower decoherence and thus improved circuit fidelity.
Figure 10. Relative number of inter-chiplet gates for SEQC normalized to the strawman baseline for devices with 10-qubit chiplets (lower is better).
Figure 10 shows the number of inter-chiplet swaps in the circuits generated by SEQC relative to the circuits generated by the baseline, with respect to the number of qubits per circuit. For essentially all tested circuits, SEQC achieves a significant reduction in the number of inter-chiplet swap gates, appearing to asymptotically converge on a roughly 4×4$\times$4 × reduction, with the greatest reduction of 4.6×4.6$\times$4.6 × at $800$ qubits / $80$ chiplets. This aligns with our model of inter- vs intra-chiplet gates, where inter-chiplet SWAPs are assumed to have 4×4$\times$4 × the error (smith_2022_chiplets) and 4×4$\times$4 × the gate duration.
Figure 11. Relative circuit depth of SEQC ormalized to the strawman baseline for devices with 10-qubit chiplets (lower is better).
Figure 12. Relative gate count of SEQC normalized to the strawman baseline for devices with 10-qubit chiplets (lower is better).
Relative circuit depth and gate count, shown in Figures 11 and 12, respectively, share similar behavior and analysis. At small circuit sizes, SEQC has higher circuit depth and gate count than the baseline. The gap shrinks as the circuit size increases, although SEQC always tends to perform worse on these metrics on average. This behavior is to be expected, since the baseline directly optimizes for circuit depth and gate count, while SEQC prioritizes other metrics, e.g., reducing inter-chiplet gates.
6.1. Discussion
The discrepancy between the circuit depth and gate count results compared to the execution time and fidelity results, respectively, are indicative of the differences between SEQC and the contemporary quantum compilers. Circuit depth and gate count represent the heuristics used in the baseline, especially SABRE (sabre_swap), and they are fundamentally unweighted metrics. In other words, circuit depth and gate count only consider the number of gates (in the critical path and in total) irrespective of each one’s unique underlying properties. Meanwhile, SEQC is able to more directly estimate execution time and fidelity, the practically relevant metrics, by constructing weighted heuristics.
With the context of our static metrics (number of inter-chiplet gates, circuit depth, and gate count), it is possible to justify the improvements in relative fidelity and relative per-shot execution time as the circuit size increases. Namely, as expected, the unweighted heuristics in SABRE tend to under-represent the error and gate delay contribution of inter-chiplet gates. This effect gets exacerbated for wider or deeper circuits, which increase the number of opportunities for cross-chiplet interactions.
Overall, we observe that SEQC yields circuits with more desirable characteristics when executing on modular quantum architectures, while at the same time it achieves faster compilation times. Also, as the number of qubits increase, the performance improvement of SEQC widens compared to the baseline, especially for the most important metrics of fidelity, compiled circuit execution time, and number of inter-chiplet gates. These trends hold promise in allowing for faster and better compilation for future large-scale modular quantum devices than we can achieve with today’s frameworks.
7. Related Work
Even though quantum computing is still nascent, there is a significant body of work on techniques for compiling quantum circuits for NISQ processors. Here, we focus on the work that best relates to our contributions.
MECH (mech) sacrifices some ancillary qubit resources to build an ”adjustable multi-entry communication highway” through software. By entangling ancillary qubits and distributing them throughout the chiplet architecture’s topology, the work virtualizes those qubits into fast network-on-chip routing channels without hardware modifications. Future work to extend SEQC with MECH’s methodology may enable combining MECH’s techniques with SEQC’s allocation procedures to gain the benefits of each.
Noting the scaling challenges and fidelity issues associated with large, monolithic quantum machines, CutQC (cutqc) establishes a hybrid computing paradigm which enables smaller-scale quantum computers to evaluate quantum circuits they would have otherwise had too few qubits to execute. To establish said paradigm, the work introduces a novel circuit-cutting technique for partitioning larger quantum circuits into small subcircuits. Classical post-processing is performed on the results of the subcircuit evaluations, merging them to find the result that would be generated by the uncut circuit. Circuit cutting as performed in CutQC served as an inspiration for SEQC’s stratification step, but unlike CutQC, SEQC does not perform any classical “stitching” of the results afterwards. It is interesting to note that our first version of SEQC was formulating the problem using constrained optimization as well, but the slowdown of the stratification step was unacceptable and impractical for large quantum circuits, and thus was abandoned in favor of the heuristics we currently employ in our compiler.
Bandic et al.(bandic2024profiling) also note the benefits of modularity. Their study is motivated by application-specific optimization – they seek techniques that allow a quantum circuit’s structure to influence hardware architecture. In their work, the relationship between circuit structure and mapping efficiency is defined, and this information optimizes the architecture of single- and multi-core quantum processors.
Piveteau and Sutter (circuit_knitting) investigate the benefits of using quasiprobability simulation to allow subcircuits cut from larger quantum circuits to perform their quantum operations independently, simulating halo qubit behavior where the subcircuits connect. When augmented with classical communication between circuits to reduce the simulation overhead incurred by sampling, they demonstrate that resulting circuits reach fixed accuracy with less simulation time. Unlike SEQC that targets quantum compilation, this technique has been developed to partition large quantum circuits into subcircuits that fit on smaller devices, at the cost of a simulation overhead to classically combine them.
Lin et al. (let_every_qubit) recognize that each 2-qubit gate in a quantum circuit has a dynamic behavioral profile which depends on manufacturing and environmental factors, and present a compilation technique which uses each 2-qubit set’s calibration data to determine potentially nonstandard 2Q gate operations which are much more ideal for that set to perform than their standard counterparts—essentially, choosing a more favorable basis. By characterizing these nonstandard gates, using their unitaries to synthesize standard gates, and essentially allowing 2Q gates to perform standard operations along bases more compatible with their respective physical states, the technique improves fidelity. This technique is orthogonal to SEQC and could work synergistically with it.
By mapping circuit partitioning to a minimum cut problem, Bandic et al. (bandic2023mapping) use the QUBO model with quantum solvers to generate a set of cuts intended to minimize swaps between circuit partitions. QUBO employs quantum optimization algorithms on quantum annealers to accelerate the solution process and remain practical, as the search space for transmission qubits is $2^{n}$. SEQC, in contrast, does not require acceleration through quantum means, but at the cost of not guaranteeing an optimal solution (it is rather a “best-effort” approach that works well in practice).
Through discussion of scalability issues impeding the development of larger monolithic quantum circuits, Ovide et al. (ovide2023mapping) motivate the necessity of mapping quantum circuits to multi-core architectures and present a commonly agreed-upon likely computing paradigm which leverages both quantum and classical methods to foster inter-chip communication, circumventing the scalability challenges inherent to monolithic quantum processors. While this work advocates for a two-level, hierarchical approach for quantum circuit mapping, it leaves the actual problem of qubit placement and then routing unsolved. SEQC provides an algorithm and implementation, addressing both of these challenges.
Using iterative optimization via the Hungarian algorithm, Hungarian Qubit Assignment (HQA) (escofet2023hungarian) seeks to optimize the qubit-to-core mapping problem. Because the Hungarian algorithm is here extended with the ability to adjust a qubit’s cost matrix according to behavior multiple timeslices ahead, the model can leverage lookahead to optimize further the number of inter-core swaps. However, HQA operates under the assumption that all inter-chiplet communications have the same cost, meaning chiplets must be fully connected. SEQC, on the other hand, is able to work with any arrangement of chiplets and inter-chiplet links (as long as there is a path between any two chiplets), taking into account the distance between pairs of chiplets to generate improved circuits.
Iterating on their previous work on Hungarian Qubit Assignment, Escofet et al. (escofet2024revisiting) develop a characterization technique for arbitrary quantum circuits which yields theoretical upper and lower optimization bounds. This is significant, since it enables optimality assessment for solutions to the qubit mapping problem.
Pastor et al. (pastor2024circuit) leverage deep reinforcement learning to partition quantum circuits for execution across multiple processing cores. Using a DRL model enhanced with policy-masking mechanisms and a lookahead-capable observation array, the work finds valid qubit-circuit mappings for each of timeslice while prioritizing the reduction of inter-core swaps. Notably, it only considers inter-core swaps for optimization, deeming intra-core single-and-multiple qubit operations so inexpensive that their optimization should serve as fodder for inter-core swap reductions. This is sensible, but leaves optimization potential on the table, since it only seeks to generate valid mappings within each timeslice. Additionally, similar to HQA, this technique treats all inter-chiplet communications equally, further missing optimization opportunities with non-fully connected chiplet topologies. SEQC, in contract, is cognizant of the differing costs of moving information between each pair of chiplets, and optimizes accordingly.
To solve the qubit allocation (mapping) problem, Liu et al. (liu_2023_tackling) develop an algorithm that combines partitioning, permutation-aware synthesis, and a modified version of SABRE. In a bottom-up approach, this technique partitions a logical circuit into blocks, performs permutation-aware synthesis on each block in parallel (solving for all possible qubit topologies), and finally uses an augmented SABRE algorithm to perform initial layout and routing between blocks. In contrast, SEQC takes a top-down approach, where SWAP permutations between chiplets are settled prior to any compilation tasks performed within each block. As a consequence, we only need to generate solutions for one qubit layout per block, rather than synthesizing all possible block permutations.
Parallelizing SEQC’s recurrent elaboration step was inspired by traditional classical compilation techniques, which has a decades-long history and is now commonplace (Mickunas_1978_ParallelCompilation; Gross_1989_ParallelCompilation). Besides traditional compilation, though, we can also draw parallels between our work and prior research on data movement in the classical space. Data movement is widely understood to be the primary performance bottleneck in the classical architecture space. In the last two decades, work has proliferated on dataflow and spatial architectures, much of which focuses on mapping problems. Consider neural network architectures, for example, which seek to map their physical topologies as closely as possible to problem topologies. Eyeriss (Eyeriss_2016_Chen) approaches the dataflow mapping problem by proposing a Row-Stationary architecture, and then optimizing the mapping of problem to processing element such that elements which are spatially local in the problem are spatially local in hardware during execution. This is not unlike qubit swap optimization, and in fact is performed for many of the same reasons—like inter-chiplet swaps, data movements between more distant classical programming elements in a dataflow architecture are simply more expensive, and minimizing their frequency yields performance improvements.
8. Conclusion
Physical, technological and economic considerations make the prospect of a single “monolithic” large-scale quantum processor very expensive or entirely unrealistic by current standards. These constraints force future quantum systems to adopt a physically distributed architecture, based on modular quantum chiplets. As such, modular quantum processors already dominate the roadmaps of several major quantum companies.
However, the complexity of chiplet-based quantum devices, coupled with their growing size, presents an imminent scalability challenge for quantum compilation. Existing qubit allocation methods are often unable to contend with inter-chiplet links, which don’t necessary support a universal basis gate set, and existing methods of logical-to-physical qubit placement, routing, unitary synthesis, and quantum circuit optimization, are typically not designed for qubit links of wildly varying levels of duration or fidelity.
In this work, we first modify stock contemporary compilers to become aware of hardware modularity, thereby allowing them to correctly compile circuits for modular architectures with limited cross-chiplet gate support. We then propose SEQC, a Stratify-Elaborate Quantum Compiler for modular architectures. SEQC performs compilation in two stages, with the first stage (stratification, or chiplet splitting) performed only once for a given architecture, and the second stage (elaboration, or chiplet compilation) performed in parallel for each chiplet. Unlike contemporary quantum compilers that have to compile a quantum circuit anew before each execution, only the elaboration stage of SEQC needs to be performed recurrently before each execution. As its computational complexity is a function of the quantum chiplet size, not the entire quantum processor size, and chiplet sizes grow slowly, SEQC’s recurrent compilation time is largely unaffected by the relentless growth of qubit counts in future quantum processors.
We design and implement several novel methods for qubit placement, qubit routing, and circuit optimization in SEQC, with both hardware modularity and compilation parallelization in mind. Owning to these techniques and its parallelization strategy, SEQC compiles circuits that exhibit up to a 36% increase in circuit fidelity and up to 1.92×1.92$\times$1.92 × lower execution time, while consistently achieving 24×2-4$\times$2 - 4 × faster compilation time over a chiplet-aware Qiskit baseline.
[…]