### v4

# ab,ab.braket

Abstract.
[…]
Quantum computing, quantum compilation, modular architectures.
copyright: nonejournalyear: 2024doi: XXXXXXX.XXXXXXX
1. Introduction
Classical computer systems, through the decades of their existence, have become increasingly distributed. Physical, technological, and economic constraints have prevented single “monolithic” systems from scaling to the point that they could meet the demand for high performance and yet remain practical. Today, distributed architectures are prevalent, and their latest renditions, from cloud computing to chiplet-based digital processors, are ubiquitous.
We postulate that quantum computing is on a similar path. While improvements in superconducting quantum hardware have led to the debut of processors with $1000+$ qubits (ibm_condor), practical implementations of quantum computation will require millions of physical qubits (paler2019really). The number of qubits required by a quantum algorithm, the depth of the algorithm (i.e., execution time or gate count on the critical path), the auxiliary and syndrome qubits required to reach sufficiently low error rates, and the size of qubits with all control hardware, all suggest that many more qubits are required than are likely to fit in a single die (vanMeter_2010_DQC). The overhead for error correction, cooling, dilution, control systems, I/O lines, challenges associated with verification and adequate chip thermalization, and realistic resources such as lossy waveguides, limited qubit fabrication yields as qubit capacity grows (smith_2022_chiplets), and finite chip sizes, make the prospect of a single “monolithic” quantum processor very expensive or entirely unrealistic by current standards. These constraints force large-scale quantum systems to adopt a physically distributed architecture.
We are already observing signs of this shift. Recent developments in quantum chip linking, including flip-chip architectures (gold2021entanglement) and low-loss coaxial cables (niu2023low), suggest that modular designs are the most viable for scaling (smith_2022_chiplets). Many contemporary or upcoming leading quantum systems adopt chiplet-based modular quantum processors, for example using high-bandwidth quantum links between nearest-neighbor quantum chiplets (e.g., carrier-chip couplers in Rigetti Aspen-M (Rigetti_AspenM; gold2021entanglement), or m-couplers in IBM Crossbill (gambetta_2024_IBMcouplers; Mandelbaum_2024_IBMCouplers)), or lower-fidelity, lower-bandwidth, but longer-distance flexible coupling of discrete chips (e.g., l-couplers in IBM Flamingo (gambetta_2024_IBMcouplers; Mandelbaum_2024_IBMCouplers)), or a combination of the above (e.g., c-, m-, and l-couplers in IBM Starling (Mandelbaum_2024_IBMCouplers)), or multi-chip connectivity through tunable couplers and routing chips in Rigetti Cepheus (Rigetti_Cepheus; Field_2024_RigettiModular). Most major companies have set their sights at modular designs to meet quantum scaling targets in practical ways, and modular quantum processors dominate their latest roadmaps (IBM_2025_FTQC; IonQ_2024_Roadmap; Rigetti_2024_Roadmap).
At the same time, no single qubit technology simultaneously offers long coherence, fast gates, high connectivity, and scalable, mature fabrication. Superconducting qubits offer very fast gate latencies and mature fabrication, but low coherence times and mostly local connectivity. Trapped ions have excellent coherence but slow speeds, and all-to-all connectivity but only for the few ions located at the C-cut sapphire window. Neutral atoms have excellent scalability and arrangeable connectivity, but do not reach the gate speeds and manufacturing robustness of superconducting qubits, or the coherence times of trapped ions. Hybrid, modular quantum architectures, comprising QPUs based on different qubit modalities, provide a path to combine the complementary strengths of each platform, enabling performance impossible for any single modality. For example, superconducting qubits can serve as fast computational QPUs, while ions can provide memory QPUs, and Rydberg atoms can provide ancillae and distillation functionality. Such heterogeneity mirrors the classical transition from monolithic CPUs to systems composed of CPUs, GPUs, NPUs, and specialized accelerators. As such, hybrid modular quantum systems are increasingly recognized as a pragmatic design path for long-term universal quantum computers (DARPA_HARQ; HQAN; hetarch; Xiang_superconducting_hybrid_2013). Efforts to realize hybrid quantum systems have predominantly leveraged quantum state transfer, i.e., SWAPs, between heterogeneous qubit technologies (Kurizki_hybrid_2015; Clerk2020; Zhou2014; PhysRevLett.108.130504). While universal quantum operations, e.g., CNOT or CZ operations, that interface two different quantum technologies have been demonstrated (PhysRevLett.114.080501; Xiang_superconducting_hybrid_2013; Cia_2024; Daniilidis_2013; Yu2016), these often have to compromise on gate fidelity, and have yet to be realized for all combinations of common qubit technologies, e.g, between superconducting and trapped-ion qubits. Thus, hybrid quantum systems may exhibit interconnect heterogeneity, where inter-module links may not be able to support universal gate sets, while intra-module couplers do.
Unfortunately, current compilation infrastructure does not support quantum interconnect heterogeneity. The mapping of a quantum program’s logical quantum gates into the native gates supported by the underlying quantum processor, and the mapping of logical qubits to the physical qubits of the quantum hardware, largely determine the number and type of native quantum gates required for the computation, the circuit depth, and its execution time—long programs may not complete successfully as qubits decohere. Different mappings can result in vastly different compiled quantum circuits, with diverse characteristics along all these axes. These details of efficiency can make or break an algorithm in today’s noisy intermediate-scale quantum systems (NISQ) era, so quantum compilers aggressively optimize for all of them. To make matters more challenging, each qubit, coupler, and gate have diverse error profiles that are highly variable both spatially and temporally (Tannu_2019_ErrorVariability; Murali_2019_NoiseMapping; dasgupta2021stability). Unlike classical compilation that is done only once for a given architecture, quantum programs must be recompiled every time before execution [^1] The term “execution” here includes all shots required to produce a final result for the user. Different executions of the same circuit can span qubit calibration windows., as the ideal physical qubits to target change between runs. So, quantum compilers today are faced with the daunting task of optimizing quantum programs across multiple dimensions with often conflicting demands, in a continuously changing environment. Coupled with the hardness of synthesizing unitaries (which is exponential in the number of qubits) and the need for quick and frequent compilation, modern quantum compilers have no option other than to rely on heuristics to perform the task. These heuristics result in a compilation complexity that is quadratic to the number of qubits in the quantum processor (experimentally observed in Figure 11(a)). Hardware modularity, though, also presents an opportunity. Inspired by classical compilation, in this paper we leverage hardware modularity to achieve compilation modularity.
Compilation in classical systems is typically performed independently and in parallel for each source file, producing one object file per source. The individual object files are then linked together to construct the executable. We propose a compilation framework for modular quantum processors that works in a similar fashion: it stratifies, i.e., splits, the source quantum circuit into subcircuits small enough to fit in each module, and maps subcircuits to modules, and then in parallel elaborates each subcircuit and compiles it for its target module. This Stratify-Elaborate Quantum Compiler (SEQC) stratifies a source program only once for a given module architecture, and performs only the elaboration step recurrently before each execution.
As the qubit capacity $n$ of quantum processors grows exponentially, the $O(n^{2})$ compilation latency of contemporary compilers rises even faster. SEQC replaces the recurrent $Oa(n^{2})$ compilation step with several parallel $O(k^{2})$ elaboration steps for $k$-qubit modules. The stratification step is $O(n^{2})$, so the end-to-end complexity remains the same, but stratification is performed only once. The recurrent compilation in SEQC is only $O(k^{2})$, and as the number of qubits $k$ per module is expected to remain relatively stable, or grow at a much slower pace in the foreseeable future (IBM_2025_FTQC), SEQC’s recurrent compilation latency will barely grow as processors scale.
Additionally, as SEQC is cognizant of hardware modularity, it can stratify the source quantum circuit into subcircuits and map them to modules while minimizing cross-module communication. As we show in this paper, SEQC produces circuits with shorter execution times and significantly fewer inter-module gates compared to today’s stock compilers, leading to higher execution fidelity. More importantly, as the number of qubits in a processor grows, SEQC achieves even higher performance in these figures of merit.
In summary, the contributions of this paper are as follows:
  • We make stock compilers aware of modularity through a “peephole” correction pass, thereby allowing them to correctly compile circuits for modular architectures with limited cross-module gate support.
  • We design and implement SEQC, a Stratify-Elaborate Quantum Compiler for modular architectures. SEQC performs compilation in two stages, with the first stage (stratification, or module splitting) performed only once for a given architecture, and the second stage (elaboration, or module compilation) performed in parallel for each module. Only this second stage needs to be performed before each execution, and thus SEQC’s compilation time is largely unaffected by the growth of qubit counts in future quantum processors.
  • We design and implement in SEQC several novel methods for qubit placement, qubit routing, and circuit optimization. In particular, we classify swaps into symbiotic, commensalistic, and parasitic, based on how swap placement affects inter-module gates, and develop algorithms to promote symbiotic and avoid parasitic swap placement; we develop qubit “pinning” techniques to allow elaboration without disrupting module connectivity; and we adapt Simulated Annealing and Spatio-Temporal Aware Qubit Allocation (STA) techniques to aid in circuit stratification.
  • We evaluate SEQC on a modular quantum system consisting of small-scale (10-qubit heavy-hex) modules, and find that it compiles circuits with $9.3\%$ higher circuit fidelity on average (up to $49.99\%$), while consistently achieving faster compilation times ($3.27\times$ on average, up to $6.74\times$) compared to a module-unaware Qiskit baseline. When evaluated on a system with large-scale (120-qubit grid topology) modules, SEQC achieves $23.3\%$ higher fidelity on average ($63.36\%$ max), while compiling $1.34\times$ faster on average (up to $3.37\times$). More importantly, SEQC outperforms Qiskit even when the latter is afforded significant allowances (no universal gate restrictions, and much lower inter-module gate errors).
2. Background and Motivation
Compilation is often divided into distinct qubit allocation (layout and routing), basis translation, and optimization stages.
Qubit Allocation
Qubit allocation addresses the topology constraints of quantum devices, where physical qubits on a device have limited connectivity to other qubits on the device. It is critical that after this stage, which often requires the modification or addition of circuit gates, the resulting optimized, technology-dependent circuit is functionally equivalent to the original, technology-independent algorithm (smith2019quantum). As qubit allocation is NP-complete (qubit_allocation), it is typically divided into two substages, layout and routing, to make it more tractable.
Logical-to-physical qubit layout, also known as qubit placement, involves deriving an injective map from the virtual qubits of the quantum circuit to the physical qubits of the quantum device (qubit_routing; sabre_swap), cognizant of its limited connectivity. In other words, the qubit layout stage decides the initial position of logical qubits on the device.
Qubit routing operates from an initial qubit layout and inserts SWAP operations to align a quantum circuit’s gates to a given qubit topology. Consider, for example, the state-of-the-art in qubit allocation: the SABRE heuristic search algorithm (sabre_swap). In SABRE, the distinction of the layout and routing stages motivates an iterative method for solving qubit allocation: an initial layout is used to perform qubit routing, which, in turn, is used to generate a new initial layout. Because qubit routing remains NP-hard (qubit_routing), the problem is often simplified by assuming qubit links are roughly equivalent, particularly with respect to average fidelity and two-qubit gate duration. This assumption can manifest implicitly when a given device topology is represented as an unweighted, undirected graph. For monolithic architectures, one could debate the marginal tradeoff between performance and solution quality. However, for modular architectures, the cross-module connections are expected to be far worse compared to intra-module connections (smith_2022_chiplets; PhysRevLett.114.080501; Zhou2014; PhysRevLett.108.130504; Yu2016). Thus, as we will show, the current simplified model applied on modular architectures leads to worse solutions than a more physically accurate one.
(a) Example 2-module
architecture
(b) Input circuit from
qubit allocation
(c) Insertion of
perfect-shuffle SWAPs
(d) Insertion of return SWAPs
Figure 1. Peephole correction for inter-module gates that are unimplementable on hardware (i.e., non-SWAP, shown in red).
Basis Translation
Basis translation converts the gates of an input quantum circuit into gates that are physically supported on hardware. Basis translation can be performed efficiently given the target gate set is universal (mike_and_ike), and the quantum circuit contains a fixed, finite number of qubits (Solovay–Kitaev). Unfortunately, modular quantum architectures depend upon specialized links to connect otherwise-independent modules. In the general case, the cross-module links can only be used for data movement (such as SWAPs) and not computation, meaning these links may not support universal operations (niu2023low). Thus, compilers that depend on the expectation of universality among qubit links (qiskit_basis_translation_errors), and do not expect these heterogenous, non-universal basis gate sets, can generate incompatible outputs when targeting modular architectures.
Optimization
Quantum circuit optimization seeks to prune extraneous operations by finding opportunities to combine, eliminate, and parallelize quantum gates (optimization1). Optimization is highly time and resource intensive, with the problem scaling exponentially to the number of qubits and circuit depth (qfactor). As a result, partitioning, the division of a quantum circuit into blocks of quantum gates, is a common tactic among optimization techniques (qfactor; qgo; bqskit). One major consideration in this process is demarcation, i.e., where and when to draw the line between blocks. Namely, one forfeits all optimization opportunities along boundaries of demarcation, so they must be carefully chosen to minimize the negative impact on the output solution. Modular architectures are naturally synergistic with partitioning, since cross-module connections present a physically motivated boundary between operations.
3. Compilation for Hybrid Modular Quantum Systems
While SEQC can be applied on a modular quantum system comprising hybrid QPUs, in this paper we consider the simpler case of homogeneous modules as a starting point. The reasons behind this decision are simple. First, the primary goal of this paper is to introduce and evaluate the impact of the circuit stratification compilation stage with STA and Simulated Annealing we developed, along with qubit pinning, and symbiotic swap insertion. The presence of diverse-modality QPUs with varying capabilities would dilute these results, and make it hard to separate their effect from the effects of heterogeneity. At the same time, it would make comparisons to the existing state-of-the-art hard, as, to the best of our knowledge, there are no compilers that can target multiple modalities simultaneously for different parts of the same input circuit. Thus, in order to isolate the impact of our techniques, and facilitate comparisons with existing tools, in this paper we describe and evaluate SEQC assuming homogeneous modules of the same technology. When extending SEQC to a hybrid system, one only needs to execute different elaboration stages, one for each QPU modality, but the stratification stage and the other techniques we develop will remain largely unaffected. Thus, for the remainder of this paper, we will be focusing on homogeneous QPUs based on transmons, and we leave qubit-specific stratification optimizations, and the implementation of elaboration stages for other qubit species for future work.
Modular quantum architectures introduce new challenges to quantum compilation that impact the correctness of solutions and the speed of compilation. As discussed previously in §2, our model of modular quantum architectures considers the possibility that inter-module links can be non-universal. This is representative of hardware manufacturing constraints (niu2023low), where only a restricted set of operations, e.g., state transfers or SWAPs, may be supported over long-distance couplers (inter-module links). However, many compiler designs (including designs for modular architectures, see §7) experience errors or crashes when presented with non-universal, heterogenous, basis gate sets. For example, the Qiskit compiler (qiskit2024) is explicit in its compatibility for only universal basis gate sets (qiskit_basis_translation_errors). While it is possible to create a Qiskit Backend that supports heterogeneous basis gate sets, the compiler itself does not take them into account. Specifically, the Qiskit routing stage (zou2024lightsabrelightweightenhancedsabre) (an implementation of SABRE) will freely place gates on incompatible physical links, and there are no methods in the compiler to prevent or fix consequent errors. Moreover, compilers are called to target increasingly larger quantum devices, enabled by modularity. To address the mounting pressure of ever-growing compilation workloads, these compilers must be cognizant of classical hardware limitations and tradeoffs between solution quality and compilation latency.
To ameliorate these issues, we present a family of compilers optimized for modular quantum architectures that support module links with non-universal basis gate sets, and leverage modular compilation strategies. In Section 3.1, we propose a peephole compilation pass that can be inserted into any compiler to add support for non-universal basis gate sets. This would allow any compiler to be compatible with module-based devices without significant changes to their underlying implementations. Next, in Section 3.2, we introduce the stratify-elaborate quantum compiler (SEQC), which advocates for a hierarchical two-stage compilation that produces both better circuits for execution and faster compilation times.
3.1. Module-awareness via Peephole Correction
The design of the peephole correction pass was inspired by the “peephole” optimizations commonly used in classical compilation (dragon_book). Peephole passes operate within a nominal peephole window and substitute its contents with a logically equivalent replacement. Peephole correction (Figure 1) operates on a circuit generated by qubit allocation, where it detects when non-SWAP gates are placed on non-universal inter-module links (Figure 1(b)), and employs a cyclic qubit rotation, i.e., a perfect shuffle permutation, to move the aforementioned gates to nearby intra-module links (Figure 1(c)). Additional “return” SWAPs are then inserted to return the qubits to their initial locations (Figure 1(d)). Overall, for every inter-module gate in the original circuit generated by SABRE, the peephole pass inserts 4 additional SWAP gates — 2 inter-module and 2 intra-module. Ultimately, this degree of overhead is inevitable when the stock, unmodified compiler is unaware of module-specific device constraints.
For example, skipping the return SWAP operations would appear to save $2$ SWAPs, and halve the cost per inter-module gate. However, it would also introduce a net qubit permutation that propagates downstream in the circuit. From the perspective of modular compilation, this creates serial dependencies that prevent parallel execution of the pass. From a routing perspective, the resultant qubit permutations disrupt the post-routing layout of qubits, causing destructive routing errors that can only be resolved by an entirely new routing pass. Overall, such a hypothetical pass would result in substantially worse output circuits compared to the peephole method.
3.2. The Stratify-Elaborate Quantum Compiler
Figure 2. High-level diagram of SEQC. Elements in red reflect modifications from existing compiler designs.
We incorporate the hierarchical elements of module architectures into SEQC by adopting a two-stage procedure (Figure 2). The first, stratification stage, represents a one-time overhead to elevate a quantum circuit to the module level. The second, elaboration stage, represents a recurring cost to complete the compilation of a quantum circuit given the most up-to-date backend properties.
3.2.1. Stratification Stage
(a) Example module topology
(b) Qubit-to-subcircuit mapping
(c) Subcircuit-to-module routing
(d) Inter-module routing
Figure 3. Step-by-step example of inter-module compilation in the Stratification stage.
In this stage, we map a quantum circuit to the inter-module device topology (Figure 3(a) in our running example) by stratifying it into smaller, interconnected subcircuits. This can be accomplished with the following intermediate steps (Figures 3(b)3(d)). First, qubit-to-subcircuit mapping designates which logical qubits belong to which logical subcircuits. Then, as part of module allocation, we define a mapping of logical subcircuits to physical modules on the target device and route inter-module gates to conform with constraints of module connectivity.
Qubit-to-Subcircuit Mapping
Qubit-to-subcircuit mapping produces a stratified quantum circuit, where logical qubits (numbers in Figure 3(b)) are grouped into subcircuits (lowercase letters). An optimal qubit-to-subcircuit mapping minimizes the number of connections between subcircuits, thereby reducing the number of inter-module SWAPs needed to transfer data between modules during execution. To perform this task, we investigate two methods: Simulated Annealing and Spatio-Temporal Aware Qubit Allocation (STA) (ovide2024sta).
Simulated annealing is an efficient stochastic optimization technique shown to work well on NP-complete problems, such as place-and-route in VLSI algorithms (annealing1989). It is an iterative method to optimize a state against a given cost function. In each iteration, the state is “kicked,” or randomly perturbed, and the new state is greedily accepted if there is a reduction in cost. However, the algorithm may randomly also accept worse solutions to escape from local minima wells. The runtime of simulated annealing is statically determined using constant tuning parameters, thus the overall complexity is determined by the complexity of the cost and kick functions, which are typically polynomial. Additionally, since each run of a simulated annealing is randomized and independent, it is possible to run multiple trials in parallel with different initial random seeds, then pick the best result. This improves the chances of finding a (near-)optimal solution. Our simulated annealing method (Algorithm 1) starts with an initial qubit-to-subcircuit mapping, which we minimize against a cost function, defined as the number of inter-module gates given a certain layout. By minimizing the number of inter-module gates, we implicitly reduce the inter-module connectivity. The kick function then swaps random qubits in the layout, ensuring that each SWAP is between qubits in different subcircuits. In doing so, we hope to move strongly interacting qubits close together, minimizing the number of cross-module gates. As input, our algorithm takes the input circuit $Circ$, the module count $N_{C}$, the module size $N_{Q}$, the starting temperature $T_{0}$, the number of changes to do per degree of temperature $\alpha$, and the cooling factor $\beta$. The last three are constant tuning parameters across runs.
Algorithm 1 Qubit-to-Subcircuit Mapping with Simulated Annealing
1: procedure Cost($Circ,L$)
2:   $C\leftarrow 0$ $\triangleright$ Count
3:   for $g\in\textsc{TwoQubitGates}(Circ)$ do
4:     $q_{0},q_{1}\leftarrow\textsc{Qubits}(g)$
5:     $s_{0},s_{1}\leftarrow L[q_{0}],L[q_{1}]$
6:     Increment $C$ if $s_{0}\neq s_{1}$
7:   return $C$
8: procedure Kick($D,N_{C},N_{Q},L$)
9:   $L^{\prime}\leftarrow\textsc{Copy}(L)$
10:   for $i=1,\dots,D$ do
11:     $s_{0}\leftarrow\textsc{RandChoice}(\{1,\dots,N_{C}\})$
12:     $s_{1}\leftarrow\textsc{RandChoice}(\{1,\dots,N_{C}\}-\{s_{0}\})$
13:     $q_{0},q_{1}\leftarrow\textsc{RandChoice}(L^{\prime}[s_{0}]),\textsc{RandChoice}(L^{\prime}[s_{1}])$
14:     $L^{\prime}[s_{0}],L^{\prime}[s_{1}]\leftarrow(L^{\prime}[s_{0}]\cup\{q_{1}\})-\{q_{0}\},(L^{\prime}[s_{1}]\cup\{q_{0}\})-\{q_{1}\}$ $\triangleright$ Swap $q_{0}$ and $q_{1}$
15:   return $L^{\prime}$
16: procedure Q2SMappingAN($Circ,N_{C},N_{Q},T_{0},\alpha,\beta$)
17:   $L\leftarrow$ Initial Layout $\triangleright$ Current Solution
18:   $R\leftarrow L$ $\triangleright$ Best Solution
19:   $T\leftarrow T_{0}$ $\triangleright$ Annealing “Temperature”
20:   while $T>1$ do
21:     $D\leftarrow\textsc{Ceil}(T\times\alpha)$ $\triangleright$ Number of changes to make
22:     $L^{\prime}\leftarrow\textsc{Kick}(D,N_{C},N_{Q},L)$
23:     $\Delta\leftarrow\textsc{Cost}(Circ,L^{\prime})-\textsc{Cost}(Circ,L)$
24:     $k\leftarrow\textsc{Rand}(0,1)$ $\triangleright$ Random number in $0,1)\subset\mathbb{R}$
25:     $L\leftarrow L^{\prime}$ if $(\Delta<0)$ or $(k>\exp(\Delta/T))$
26:     $R\leftarrow L^{\prime}$ if $(\textsc{Cost}(Circ,L^{\prime})<\textsc{Cost}(Circ,R))$
27:     $T\leftarrow T\times\beta$ $\triangleright$ Decrease temperature
28:   return $R$
The STA algorithm is a qubit allocation algorithm designed for trapped-ion quantum systems. These devices often feature modular designs, where one device is composed of “traps” arranged in a (typically) line or ring topology, where each trap contains multiple qubits (ions). Qubit allocation algorithms for modular trapped-ion systems are incentivized to minimize the number of operations between modules due to associated qubit shuttling and relocation overheads (ovide2024sta). When compiling for modular transmon architectures, we are similarly incentivized to minimize the number of inter-module interactions. Thus, we adapted a component of the STA algorithm (Algorithm 2 in Ovide et al. (ovide2024sta)) that assigns qubits to ion traps to perform qubit-to-subcircuit mapping and generate an initial subcircuit-to-module mapping (Algorithm 2).
The original STA algorithm prioritizes mapping qubits that frequently interact with other qubits (quantified by $R$ in Algorithm 2), or have a large degree of direct or indirect interactions (quantified heuristically as the depth-weighted sum $T$ of interactions between every pair of qubits). These are used to derive a sequence of qubit pairs, from which STA constructs a qubit mapping. Namely, for each iteration of the STA algorithm, a pair of qubits $(q_{a},q_{b})$ is used to insert at least one qubit into the mapping, based on the shortest topological distance between physical qubits. If one of the qubits in the pair is already mapped, say $q_{a}$, then STA also recursively maps any qubits that are more strongly connected to $q_{a}$ than $q_{b}$ is. This is done in order to keep strongly connected qubits physically close together.
Algorithm 2 Qubit-to-Subcircuit Mapping with STA
1: procedure ComputeRatios($C$)
2:   $R\leftarrow$ empty map
3:   for $q\in\textsc{Qubits}(C)$ do
4:     $R[q]\leftarrow\frac{\textsc{InteractionCnt}(q,C)}{\textsc{NumQubits}(C)}$ $\triangleright$ fraction of qubits $q$ interacts with
5:   return $R$
6: procedure ComputeST($C,S$)
7:   $T\leftarrow$ empty map
8:   for $(q_{1},q_{2})\in\textsc{AllPairs}(\textsc{Qubits}(C))$ do
9:     for $(i,s)\in\textsc{enumerate}(S)$ do
10:      Add $2^{-i}$ to $T[(q_{1},q_{2})]$ if $q_{1}$ and $q_{2}$ interact in $s$
11:   return $T$
12: procedure MapQubit($q_{1},T,R,L$)
13:   $P\leftarrow$ first pair in $T$ containing $q_{1}$; $q_{2}\leftarrow$ other qubit in $P$
14:   if $\textsc{Index}(q_{2})<\textsc{Index}(q_{1})$ then
15:     $\textsc{MapQubit}(q_{2},T)$ $\triangleright$ Map $q_{2}$ if it comes first
16:   if $q_{1}\in R$ or $q_{2}\in R$ then
17:     $\textsc{PlaceQubits}(q_{1},q_{2},L)$
18:     $R.\textsc{remove}(q_{1},q_{2})$; $T.\textsc{remove}(P)$ $\triangleright$ Remove placed qubits
19: procedure Q2SMappingSTA($C$) $\triangleright$ $C$: Quantum Circuit
20:   $S\leftarrow\textsc{Layers}(C)$; $L\leftarrow$ initial layout
21:   $R\leftarrow\textsc{ComputeRatios}(C)$; $T\leftarrow\textsc{ComputeST}(S)$
22:   while $R$ is not empty do
23:     $\textsc{MapQubit}(R[0],T,R,L)$
24:   return $L$
Module Allocation
Algorithm 3 Inter-Module Routing
1: procedure FindSwaps($Q,E$)
2:   return $\textsc{FindSymbioticSwaps}(Q,E)$ if non-empty
3:   return $\textsc{FindCommensalisticSwaps}(Q,E)$ if non-empty
4:   return $\textsc{FindParasiticSwaps}(Q,E)$
5: procedure ValidGates($Q,L$)
6:   Initialize $V_{virt}$ and $V_{phys}$ to empty sets
7:   for $g\in Q$ do
8:     $H\leftarrow\{L[q]|q\in\textsc{Qubits}(g)\}$ $\triangleright$ Get module info
9:     if $|H|=1$ then $\triangleright$ All qubits on same module
10:      Add $g$ to $V_{Virt}$ and $(g,H)$ to $V_{phys}$
11:   return $V_{virt},V_{phys}$
12: procedure Step($Q,R,L,D$) $\triangleright$ Advances the “wavefront”
13:   $V_{virt},V_{phys}\leftarrow\textsc{ValidGates}(Q,L)$
14:   $R\leftarrow R+V_{phys}$
15:   for $g\in V_{virt}$ do
16:     Remove $g$ from $Q$
17:     for $c\in\textsc{Children}(g,D)$ do
18:      Add $c$ to $Q$ if $p\in R$ for all $p\in\textsc{Parents}(c,D)$
19: procedure Heuristic($S,R,Circ$)
20:   $L^{\prime}\leftarrow\textsc{Swap}(\textsc{Copy}(L),S)$; $C\leftarrow 0$
21:   for $g\in\textsc{TwoQubitGates}(Circ)$ do
22:     if $g\notin R$ then
23:      $q_{0},q_{1}\leftarrow\textsc{Qubits}(g)$
24:      $s_{0},s_{1}\leftarrow L[q_{0}],L[q_{1}]$
25:      Increment $C$ if $s_{0}\neq s_{1}$
26:   return $C$
27: procedure InterModuleRouting($Circ,L,T$) $\triangleright$ $L$ is from Algorithm 2, $T$ is the module topology.
28:   $D\leftarrow\textsc{CircuitToDAG}(Circ)$
29:   $E\leftarrow$ All-pairs distances from module topology $T$
30:   $Q\leftarrow\textsc{FirstLayer}(DAG)$; $R\leftarrow$ empty list
31:   $\textsc{Step}(Q,R,L,D)$
32:   while $Q$ is not empty do
33:     $\mathcal{S}\leftarrow\textnormal{{FindSwaps}}(Q,E)$ $\triangleright$ SWAP Candidates
34:     $S\leftarrow\operatorname*{arg\,min}_{s\in\mathcal{S}}\textsc{Heuristic}(s,R,Circ)$
35:     $\textsc{Swap}(L,S)$
36:     $\textsc{Step}(Q,R,L,D)$
37:   $D^{\prime}\leftarrow$ DAG built in topological order from $R$ $\triangleright$ $R$ is ordered
[…]
Module allocation describes the process of transforming a stratified circuit to be compatible with the inter-module topology of a modular device. In Figure 2, we decompose module allocation into subcircuit-to-module mapping and inter-module routing, analogous to how qubit allocation can be decomposed into qubit layout and routing. Accordingly, subcircuit-to-module mapping (Figure 3(c)), defines an initial placement of logical subcircuits (lowercase letters) to physical modules (uppercase letters). Subsequently, inter-module routing (Figure 3(d)) inserts inter-module SWAPs to account for limited cross-module gate connectivity.
In our design, we use our qubit-to-subcircuit mapping to provide an initial subcircuit-to-module mapping. Then, we extend and modify the SABRE algorithm (sabre_swap) to apply to modular architectures, placing iterations of inter-module routing alongside randomized qubit permutations to inform a final subcircuit-to-module mapping. We did explore a design that extended other components of STA to perform subcircuit-to-module mapping on transmon module systems, but it underperformed this iterative approach, and had increased compilation time, hence it was abandoned.
Inter-module routing seeks to consolidate the qubits of (non-SWAP) inter-module gates to a single module by inserting inter-module SWAPs into the stratified circuit. For a single two-qubit, inter-module gate, we represent inter-module routing as a shortest-path problem, where the two qubits of the gate correspond to module endpoints in the unweighted inter-module connectivity graph. From the shortest path, we can derive an optimal sequence of SWAPs for routing, where the unweighted path length, or module distance, indicates the length of the SWAP chain. However, in the context of a stratified circuit, composed of multiple inter-module gates, it becomes possible to interleave SWAP chains (i.e., efficiently reorder their execution) to reduce the depth and/or total number of inter-module SWAPs. Like qubit routing, we would expect it to be impractical to derive an optimal solution for this more complex picture of inter-module routing at compile-time. Thus, we employ a similar SWAP selection scheme as SABRE (sabre_swap), where inserted SWAPs are chosen from a pool of SWAP candidates according to a cost heuristic. For our inter-module routing algorithm, we categorize SWAPs from our candidate pool into three tiers, based on a concept of the “efficiency” of each SWAP, as shown in Algorithm 3.
An inter-module SWAP is considered symbiotic if it directly affects two inter-module gates in a “beneficial” manner. In other words, each qubit of a symbiotic SWAP corresponds to a different inter-module gate, where the insertion of the SWAP reduces the module distance score of both gates. In theory, this outcome represents the most efficient possible insertion for a single SWAP, and thus should be the highest-priority case for our greedy inter-module routing cost heuristic. Next, commensalistic SWAPs are those that do not “harm,” or increase that module distance, of any directly impacted gates. These SWAPs typically occur when an idle qubit can be used as an ancilla qubit to facilitate the routing of an inter-module gate without affecting other inter-module gates. In certain device topologies, a SWAP acting on the qubits of two inter-module gates can also exhibit commensalistic behavior. Finally, parasitic SWAPs act on the qubits of two inter-module gates, where one gate benefits while the other is harmed. Since this variety of SWAPs is actively counterproductive towards a gate, it should be, and likely will be, avoided in most circumstances. However, given a particularly onerous circuit, it may be necessary in such rare instances to break out of a deadlock. For example, clearing one gate may free physical qubits to help route others.
3.2.2. Elaboration Stage
(a) Intra-module topology
example
(b) Stratified 3-qubit subcircuit
(c) Subcircuit after intra-module layout
$q_{0},q_{1},q_{2}]\Rightarrow[Q_{2},Q_{0},Q_{1}]$
(d) Inter-module SWAP lowering
(e) Intra-module routing
(f) Basis translation and optimization
Figure 4. Step-by-step example of intra-module compilation in the Elaboration stage. Gates that cannot be implemented on hardware are shown in red.
Following the stratification stage, each module can be compiled almost completely independently from one another. More precisely, the compilation process within each module is largely identical to that of a monolithic quantum architecture. The only exception is the presence of inter-module SWAPs, which must be lowered from the module level to the qubit level. Once this assignment is made, these SWAPs must remain fixed, which we call “qubit pinning” and immutable to subsequent compilation stages. Despite this overhead, restricting the task of compilation to a narrower domain of qubits facilitates both parallelization (since each subcircuit can be processed in parallel) and problem size reduction (since there are fewer qubits to consider) of the most computationally intensive compilation stages, namely routing and optimization. To accomplish these tasks, we divide the elaboration stage into two substages: qubit allocation and parallel basis translation and optimization. Figure 4 illustrates this with an example of elaboration for a single module (Figure 4(a)) and its corresponding subcircuit (Figure 4(b)).
Qubit Allocation
We employ a modified version of SABRE to perform qubit allocation in the elaboration stage. Specifically, we parallelize the SABRE routing algorithm and add constraints to maintain correctness when applied to a stratified quantum circuit with inter-module gates. We also explored a layout pass that used portions of the STA algorithm, but the marginal benefits to solution quality were not worth the $5-6\times$ slower elaboration time.
The placement of inter-module gates is a non-parallelizable task (due to cross-module interactions) that strongly influences the layout and routing passes, which could otherwise be fully parallelized across modules. To account for this, we introduce a serial stage (Figure 4(d)) between the parallel layout and routing passes that greedily places gates on the nearest valid inter-module edge. We choose a greedy policy because inter-module gates have more restrictions on valid placements compared to intra-module gates, while also experiencing substantially higher error rates (smith_2022_chiplets). Once the positions of the inter-module gates are settled, they are pinned, imposing an additional constraint on qubit routing. To softly encourage this behavior in SABRE, we extend its cost heuristic to weigh the device topology by fidelity. This causes higher error links to assume higher costs, discouraging SABRE from moving the high-error inter-module gates. As a result, SEQC achieves better solutions in the common case. However, there is nothing to prevent SABRE from incorrectly modifying any inter-module SWAPs. Thus, we need to impose hard restrictions to ensure correctness. To address this we establish two disjoint gate sets: one gate set that operates strictly within a module (the traditional gate set for a monolithic processor), and another that is the only gate set permitted to operate across modules (and, specifically for our example implementation, only implements inter-module SWAPs). By restricting each gate set to its own disjoint part of the overall device topology, we ensure correctness across all cases.
Parallel Translation and Optimization
The remaining compilation stages of basis translation and circuit optimization can be easily parallelized at the module granularity (Figure 4(f)). Basis translation, in particular, is embarrassingly parallel: gates can be translated independently from each other. Meanwhile, as discussed in §2, optimization algorithms already employ partitioning and parallelization to create tractable subproblems from intractably large circuits, facilitating a smaller problem space, lower memory usage, and faster performance. In modular quantum architectures, we physically motivate partitions through the natural module boundaries, enabling us to reap the same resource and performance benefits of partitioning without compromising on solution quality.
4. Methodology
We evaluate quantum compilers based on the quality of their solutions, their speed of compilation, and their ability to scale to large circuits/devices. To experimentally profile these characteristics, we benchmark a compiler on a suite of quantum circuits, detailed in §4.1, targeting a mock module-based quantum backend, as specified in §4.2. In total, we evaluate four compilers—a reference compiler, a peephole-augmented compiler, and two versions of SEQC—on simulated backends ranging from 1 to 100 modules. All compilers are implemented in the Qiskit SDK (qiskit2024) v1.2.41.2.4 and Python $3.12.6$, and timing results are taken on a machine with a $128$-core Ampere Altra Max ARM64 processor and $256$ GB of DDR4 RAM. For our reference compiler, we use the one built into Qiskit, parameterized with optimization_level=3. Because Qiskit is not natively module-aware, we run these experiments on a backend with universal inter-module links (§4.2). To construct the peephole compiler we augment the reference Qiskit compiler by inserting our peephole pass between its routing and basis_translation stages, thus making it module-aware. Our two SEQC versions differ in the technique they use for qubit-to-subcircuit mapping. Our first SEQC implementation leverages simulated annealing, initialized with a starting temperature $T_{0}=200$, a rate of change of $0.005T$ per round, $\frac{T}{50}$ permutations per round, and a total of $10$ trials per CPU core. Meanwhile, our second SEQC implementation leverages STA for subcircuit-to-qubit mapping and subcircuit-to-module mapping.
We aggregate results in a manner similar to SPEC benchmarks (spec2017), by taking the geometric mean of the relative performance ratio against a baseline technique across a benchmark suite of quantum circuits. The resultant score is then derived for each of the evaluated metrics on multiple backends of varying numbers of modules.
4.1. Benchmark Suite
We leverage circuits from Supermarq (supermarq) that are representative of practical, real-world applications for quantum computing, while also being feasible to compile for a large number of qubits. In particular, our suite comprises the following quantum circuits.
[…]
(a) BitCode
(b) PhaseCode
(c) GHZ
(d) VQE
(e) Hamiltonian Sim.
Figure 5. Examples of benchmark applications for 5-qubit circuits.
[…]
VQE. The variational quantum eigensolver (VQE) (vqe) is another hybrid quantum-classical variational algorithm that is often used in chemistry applications. In our experiments, we choose a $2$-layer ansatz whose parameters are again randomly generated (Figure 5(d)). The resultant circuit structure features two CNOT chains separated by single-qubit operations. The CNOT chains can be interleaved, which in conjunction with the single-qubit gates, creates many opportunities for parallelism (resembling pipelining in classical computing). Moreover, from a qubit routing perspective, the second CNOT chain forces the compiler to consider how qubit permutations in one circuit layer propagate to future layers.
Hamiltonian Simulation. Quantum devices are naturally well-suited to simulating Hamiltonians. The Supermarq benchmark selects the 1D Transverse Field Ising Model (TFIM) system to model, which is relevant for phase transitions in magnetic materials. The circuit for Hamiltonian Simulation (Figure 5(e)) resembles a slightly more complex GHZ circuit, including how it is completely serial. Compared to GHZ, this circuit should test the compiler’s ability to lookahead during routing to detect back-to-back gates on the same two-qubit pair, while also emphasizing parallelism originating from the efficient placement of SWAP gates.
4.2. Simulated Quantum Device Specifications
(a) Small-scale heavy-hex modules, on grid lattice.
(b) Large-scale modules with IBM Nighthawk-like topology.
Figure 6. Modeled topologies.
T1 (backend_specs) $20\times 10^{-6}$ seconds
T2 (backend_specs) $30\times 10^{-6}$ seconds
Frequency (backend_specs) $6\times 10^{9}$ Hz
(a) Qubit Properties
Instruction Duration (ns) Error
X (backend_specs) $25.0$ $0.109\%$
SX (backend_specs) $25.0$ $0.109\%$
$\text{R}_{\text{z}}(\phi)$ (backend_specs) $0.0$ $0.000\%$
CZ (intra-module) (backend_specs) $34.0$ $0.605\%$
CZ (inter-module) $126.0$ $2.420\%$
SWAP (inter-module) $702.4$ $10.23\%$
Reset (backend_specs) $500.0$ $0.186\%$
Measure (backend_specs) $500.0$ $0.196\%$
(b) Instruction Properties
Universal gate set only. Non-universal gate set only.
Table 1. Specifications of Simulated Modular Backend.
We model two types of simulated devices differing in their inter-module topology and module size. To isolate the impact of architecture modularity, we evaluated on symmetrical modules based on a transmon qubit technology. One variant, denoted the heavy-hex backend (Figure 6(a)), is generated to conform to the modular architecture proposed in Smith et al. (smith_2022_chiplets), and features a 10-qubit, heavy-hexagon topology (heavy_hex). The other, denoted the grid backend (Figure 6(b)), uses modules with a 120-qubit grid topology, based on the upcoming IBM Nighthawk processor (IBM_2025_FTQC; Nighthawk_topology). Both simulated devices use a grid lattice topology of symmetric modules. The hardware specifications of the qubits and intra-module connections (detailed in Table 1) are sourced from Acharya et al. (backend_specs). To generate the expected fidelity of the inter-module SWAPs, we perform a simulated random benchmarking trial and calculate the error rate of a SWAP gate with these backend specifications. Following Smith et al. (smith_2022_chiplets), we conservatively model inter-module SWAPs with $4\times$ the error rate and $4\times$ the duration of an intra-module SWAP. Finally, note that for each valid module count, there may exist multiple valid module topologies. In this work, we generate our backends to use the “most-square” topology for a given module count. For example, a $12$-module machine would be a $3\times 4$ grid.
5. Experimental Results
We evaluate the Qiskit reference compiler, the peephole-augmented Qiskit (PA-Qiskit) compiler, SEQC using simulated annealing (SEQC-AN), and SEQC using STA (SEQC-STA) on metrics relating to the quality of the compiled circuits, on both the 10-qubit heavy-hex backend, and the 120-qubit grid backend. The Qiskit-based and SEQC compilers prioritize different heuristics in compilation. We examine how well each compiler is able to minimize its preferred heuristics, and how well that translates to practical solution quality. To demonstrate scalability, we analyze compilation time, specifically the recurring and non-recurring compilation times for repeating compilations of a single application, as well as memory requirements for the compilation. Relative metrics are normalized over the Qiskit reference compiler.
Number of inter-module gates
The primary heuristic we minimize in SEQC is the number of inter-module gates in a quantum circuit. This is because we expect that the inter-module gates of a modular architecture contribute the most to error and execution time relative to other gates (at least by a factor of $4\times$ (smith_2022_chiplets)). As a result, the number of inter-module gates should represent a strong proxy metric for the quality of a compiler solution, and thus minimizing the number of inter-module gates should result in higher-fidelity, faster-executing quantum circuits. Accordingly, SEQC consistently produces results that have significantly fewer inter-module gates compared to Qiskit-based compilers (Figures 7(a) and 8(a)). For example, compilation with SEQC-STA results in $82.43\%$ fewer inter-module gates on average compared to Qiskit (up to $92.94\%$). on the heavy-hex backend. Meanwhile, compilation with SEQC-STA results in $96.31\%$ fewer inter-module gates on average compared to Qiskit (up to $98.32\%$) on the grid backend. Compared to PA-Qiskit, compilation with SEQC-STA results in $64.92\%$ fewer inter-module gates on average, up to $78.62\%$, on the heavy-hex backend. Compared to PA-Qiskit, compilation with SEQC-STA results in $94.31\%$ fewer inter-module gates on average, up to $97.83\%$ on the grid backend.
It is worth mentioning that the inter-module gates in the universal backend would be CZ gates, while the non-universal backend would be SWAP gates. For every inter-module (CZ) gate in Qiskit, PA-Qiskit should have 2 inter-module (SWAP) gates. However, we observe PA-Qiskit produces circuits with a roughly comparable number of inter-module gates to Qiskit. In fact, as size increases, PA-Qiskit tends to result in fewer inter-module gates than Qiskit. This outcome may indicate the presence of optimizations that are removing extraneous SWAPs.
(a) Relative inter-module gate count
(lower is better).
(b) Relative estimated fidelity (ESP)
(higher is better).
Figure 7. Main circuit performance metrics: inter-module gate count and estimated fidelity, 10-qubit heavy-hex backend.
(a) Relative inter-module gate count
(lower is better).
(b) Relative estimated fidelity (ESP)
(higher is better).
Figure 8. Main circuit performance metrics: inter-module gate count and estimated fidelity, 120-qubit grid backend.
Fidelity
Practical applications of quantum computing require some guarantee of program output correctness (measurements). Thus, the highest priority for quantum circuit compilation is to maximize fidelity. We estimate fidelity through estimated success probability (ESP) (esp). Specifically, we track the per-qubit ESP of a quantum circuit and report an averaged ESP for each benchmark from the geometric mean of the circuit measurements. Figures 7(b) and 8(b) presents the relative estimated fidelity of each compiler compared to the Qiskit compiler with respect to circuit size. There is a noticeable benefit in fidelity associated with the SEQC compiler, particularly with SEQC-STA. On the heavy-hex backend, we observe an average improvement of $22.13\%$ over PA-Qiskit, and up to $62.87\%$. On the grid backend, SEQC-compiled circuits achieve $3.51\%$ higher fidelity on average ($22.50\%$ max) over PA-Qiskit.
SEQC similarly outperforms Qiskit by a wide margin. It is important to emphasize here that SEQC outperforms Qiskit even when the latter is afforded significant allowances (no gate restrictions, and much lower inter-module gate errors). The Qiskit trial backend supports universal inter-module links, in contrast to the SEQC backend, and with only a $2.42\%$ error for inter-module CZs, compared to $10.23\%$ error for SEQC’s inter-module SWAPs. All these allowances should translate to higher fidelity solutions for Qiskit. Nevertheless, even though much more constrained, SEQC-STA demonstrates an average improvement of $9.30\%$, up to a $49.99\%$, on the heavy-hex backend, and $32.30\%$ ($63.36\%$ max) on the grid backend.
(a) Relative total gate count.
(b) Relative circuit depth.
(c) Relative per-shot execution time.
Figure 9. Total gate count, circuit depth, and per-shot execution time, 10-qubit heavy-hex backend (lower is better).
(a) Relative total gate count.
(b) Relative circuit depth.
(c) Relative per-shot execution time.
Figure 10. Total gate count, circuit depth, and per-shot execution time, 120-qubit grid backend (lower is better).
Gate count and circuit depth.
Gate count and circuit depth are static, backend-agnostic metrics that can serve as indicators for a quantum circuit’s expected fidelity and/or execution time, where circuit depth is the critical path of a quantum circuit in terms of individual gates (mike_and_ike). SABRE (zou2024lightsabrelightweightenhancedsabre; sabre_swap), moreover, explicitly minimizes gate count by incorporating it into its cost functions. We observe that SEQC typically produces results with higher gate count (Figures 9(a) and 10(a)) and circuit depth (Figures 9(b) and 10(b)) compared to Qiskit and PA-Qiskit. This is due to SEQC prioritizing the reduction of the number of inter-module gates over total gate count or circuit depth. Nevertheless, SEQC-STA, in particular, is able to demonstrate competitive results in these metrics for larger circuits.
SEQC-STA has comparable or slightly better results compared to PA-Qiskit at $500$ qubits (50 modules) and above for both metrics on the heavy-hex backend. For the $810$ and $1000$ qubit ($81$ and $100$ module) trials on the heavy-hex backend, SEQC-STA achieves $14.11\%$ and $8.87\%$ gate count reduction over Qiskit, respectively, despite the latter using universal inter-module links. For the grid backend, we observe similar behavior to the heavy-hex backend, where SEQC produces increasingly competitive results as the circuit size increases. However, we did not perform any trails large enough to witness the point where SEQC-STA would begin to produce better results than the Qiskit baselines
Estimated execution time.
The per-shot execution time of a quantum circuit is determined by the critical path of gate operations weighted by gate duration. In addition to performance concerns, execution time indirectly affects fidelity on NISQ-era devices due to decoherence. In other words, it is imperative that quantum circuits execute within the decoherence time to ensure high-fidelity outcomes. Figures 9(c) and 10(c) presents the estimated per-shot execution time of our proposed compilers relative to the reference as a function of the number of qubits per circuit. As explained previously, we approximate the duration of an inter-module SWAP as $4\times$ the duration of an intra-module SWAP (smith_2022_chiplets). Under this model, we observe that the SEQC compilers consistently outperform PA-Qiskit in execution by $24.19\%$ on average ($63.98\%$ max) on the heavy-hex backend. It is notable that, despite the gate set affordances given to the Qiskit compiler, SEQC-STA achieves parity in execution time starting from $600$ qubits ($60$ modules). For $810$ qubits ($81$ modules) and above, SEQC-STA actually compiles faster-executing quantum circuits. Meanwhile, on the grid backend, SEQC-STA-compiled circuits execute up to $151.29\%$ faster over PA-Qiskit. Circuit execution time for SEQC-STA-compiled circuits is also faster than the Qiskit baseline from $480$ qubits ($4$ modules) onwards by $28.04\%$ on average, up to $54.82\%$, with SEQC-SA trailing close behind.
Compilation time
Quantum devices feature rapidly fluctuating error characteristics that can change hourly, or even on a run-to-run basis (Tannu_2019_ErrorVariability; Murali_2019_NoiseMapping; dasgupta2021stability). Thus, across the entire execution lifetime of an application (quantum circuit), regular re-compilation is necessary to generate the highest-quality circuits with the most up-to-date characteristics possible for every execution. Traditional quantum compilers, like Qiskit and PA-Qiskit, approach compilation tabula rasa, where each new compilation restarts from the original algorithmic circuit. It is challenging to speed up repeated recompilations without fundamental changes to their design, as noise-aware routing and synthesis algorithms, which leverage updated error characteristics, usually function from a device-agnostic starting point. In contrast, SEQC’s stratification stage enables stateful retention of previous compilations. It is agnostic to the ever-changing error characteristics of a quantum device, and instead only depends on the device’s static topology. Therefore, a stratified quantum circuit can be perennially reused over multiple rounds of compilation, and thus SEQC only repeats the elaboration stage.
Nevertheless, although stratification is a one-time cost for the SEQC compiler, it must be scalable in order to be practical. In Figures 11(a) and 12(a), we plot the non-recurring compilation time of our evaluated compilers. For our SEQC compilers, we plot stratification time and stratification + elaboration time. These can be compared to the total solve time for the Qiskit and PA-Qiskit compilers. We demonstrate that all compilation times obey a quadratic trajectory ($R^{2}\geq 0.9883$).
Figures 11(b) and 12(b) show the recurring compilation time. Specifically, this refers to the elaboration time of SEQC and the full compilation time for Qiskit and PA-Qiskit. Overall, SEQC compilers have a large speedup over Qiskit-based compilers, owing to the use of parallelism and modular compilation strategies. Compared to Qiskit on the heavy-hex backend, SEQC-STA achieves $3.27\times$ speedup on average (up to $6.74\times$), and compared to PA-Qiskit on the heavy-hex backend, $4.69\times$ average speedup ($6.61\times$ max). Similarly on the grid backend, SEQC-STA achieves $1.34\times$ speedup on average (up to $3.37\times$) over Qiskit, and $2.52\times$ speedup on average (up to $11.93\times$) over PA-Qiskit.
(a) Compilation time.
(b) Relative recurring compilation time.
(c) Relative peak memory utilization.
Figure 11. Compilation performance metrics: compilation time, recurring compilation time, and peak memory utilization, 10-qubit heavy-hex backend (lower is better).
(a) Compilation time.
(b) Relative recurring compilation time.
(c) Relative peak memory utilization.
Figure 12. Compilation performance metrics: compilation time, recurring compilation time, and peak memory utilization, 120-qubit grid backend (lower is better).
Memory Utilization.
Figures 11(c) and 12(c) presents the relative peak memory utilization of the evaluated compilers. Compared to Qiskit, SEQC-SA exhibits $55.68\%$ higher memory usage on average ($117.71\%$ max) on the heavy-hex backend and $23.71\%$ higher memory usage on average ($38.24\%$ max) on the grid backend. This could likely be ameliorated by reducing the number of annealing trials, at the expense of circuit quality. SEQC-STA tends to use less memory compared to the Qiskit variants, by $4.22\%$ on average ($17.47\%$ max) on the heavy-hex backend and $10.69\%$ on average ($17.51\%$ max) on the grid backend. It is possible that the process of stratification and parallelized elaboration is responsible for the conservation of memory usage, presuming the memory requirements of quantum compilation scales super-linearly with the circuit size. This points to the scalability benefits of the SEQC framework in general.
6. Discussion
These results demonstrate a flaw with gate count and circuit depth as heuristics: their lack of noise-awareness. Circuit depth and gate count only consider the number of gates (in the critical path and in total) irrespective of each one’s unique underlying properties. It is expected that two-qubit gates will be significantly more noisy than single-qubit gates, and inter-module gates should be significantly more noisy than intra-module gates (smith_2022_chiplets). Failure to account for duration and error differences between gates or types of gates in quantum compilation heuristics will inevitably result in sub-optimal solutions when the discrepancies are extreme, e.g., in modular architectures. Conversely, the hierarchical structure of the SEQC framework with the stratification and elaboration stages imparts some degree of implicit noise-awareness, where minimizing the number of inter-module gates takes precedence over intra-module/total gates. As a result, SEQC can better optimize for the practically significant metrics of circuit fidelity and execution time.
Overall, we observe that SEQC yields circuits with more desirable characteristics when executing on modular quantum architectures, while at the same time it achieves faster compilation times. Also, as the number of qubits increases, the performance improvement of SEQC widens compared to the baseline, especially for the most important metrics of fidelity, compiled circuit execution time, and number of inter-module gates. These trends hold promise in allowing for faster and better compilation for future large-scale, hybrid modular quantum devices than we can achieve with today’s frameworks.
It is also important to note that SEQC is equally applicable to NISQ as well as future quantum systems. The stratification stage can decompose a large circuit into subcircuits that will be mapped to QPUs at the logical level, and then the elaboration stage can compile for each QPU together with an error correction method for that QPU. The two stages are distinct, and stratification operates more on the logical level, while the elaboration can be tasked with lowering from the logical to the physical qubits. Applicability to NISQ systems is still important, as even with technological breakthroughs, the transition to Fault Tolerant Quantum Computing (FTQC) will be gradual, not instantaneous. The hardware that exists now will continue improving and will be widely used for years, as there is practical scientific and industrial value in NISQ, even without the need to wait for fully error-corrected systems (piccinelli_chemistry; Kim2023preFTUtility; sciadv_utility_2025; Abanin2025ConstructiveInterference). Moreover, early fault-tolerant demonstrations will run on NISQ hardware, and will heavily rely on NISQ-era methods, as only few qubits can be protected. Thus, it remains critical to design architectures and compilers that serve NISQ hardware, both to extract near-term value, and to pave the engineering path toward full FTQC.
7. Related Work
The state-of-the-art in distributed quantum computing (DQC) (autocomm; 10228915; 10.1145/3730585; 10821229; 10313645; 10.1145/3579367) requires the generation and distribution of EPR pairs as the main entanglement resource over long-distance optical links, and the deployment of emitter-emitter, emitter-scatterer, or scatterer-scatterer photon-mediated entanglement schemes. However, modular architectures have less restrictive distance constraints, and they can often support other hardware technologies, such as microwave links (PhysRevApplied.21.054063). As a result, SEQC can model inter-module allocation topologically (as SWAP chains) (LaRacuente2025modelingshortrange), similar to compilation for monolithic (single-die) quantum processors. An analogy can be drawn to classical compilation between chiplet-based classical systems and distributed classical systems, where diverse compilation strategies are required due to very different hardware technologies and distances between computation units.
Compilation for modular transmon systems shares many similar constraints with compilation for quantum charge-coupled devices (QCCDs). Namely, these devices features modules, in the form of ion traps, and support movement between traps, via ion shuttling mechanisms (ovide2024sta). However, there are some key differences between the hardware technologies that make it difficult to compare SEQC with QCCD compilers on equal footing. Notably, qubits in superconducting devices needs to be topologically adjacent to each another to support two-qubit operations. However, QCCD devices are less restrictive, and support two-qubit operations so long as qubits are in the same traps / interaction region. As a result, QCCD compilers (10.1145/3526241.3530366; Kreppel2023quantumcircuit) can employ simpler routing strategies, or forgo them entirely, compared to modular transmon compilers.
MECH (mech) entangles ancillary qubits and distributes them throughout a module to build fast network-on-chip routing channels (a ”communication highway”). MECH includes a compiler that requires, and optimizes for, the “highway” design. Unfortunately, the highway does not generalize to non-universal inter-module links. Future work to extend SEQC with MECH’s methodology may enable combining MECH’s techniques with SEQC’s allocation procedures to gain the benefits of each. Lin et al. (let_every_qubit) present a compilation technique that utilizes calibration data to determine potentially nonstandard 2Q gates that can perform standard operations along bases more compatible with their respective physical states and improve fidelity. Bandic et al.(bandic2024profiling) utilize the relationship between quantum circuit structure and mapping efficiency to optimize the architecture of monolithic and modular quantum processors, on the path to application-specific hardware optimization. Escofet et al. (escofet2024revisiting) develop a characterization technique for arbitrary quantum circuits, which provides theoretical upper/lower optimization bounds, enabling the optimality assessment of qubit mapping solutions. These techniques are orthogonal to SEQC, and could work synergistically with it.
CutQC (cutqc) partitions large quantum circuits into smaller subcircuits that fit on a device, executes them one at a time, and then merges the individual results through classical post-processing. Piveteau and Sutter (circuit_knitting) similarly allow subcircuits to execute independently, and use quasiprobability simulation to simulate inter-module link behavior where the subcircuits connect, reaching fixed accuracy with less simulation time. SEQC’s stratification stage was inspired by CutQC, but unlike these works SEQC does not perform any classical “stitching” or simulation. We initially formulated qubit-to-subcircuit mapping as a constraint optimization problem, similar to CutQC, but existing numerical solvers were impractically slow for circuits larger than 20 qubits, corroborating the experience of Baker et al. (baker2020). Instead, we opted to use the approximate, but practical, solutions we currently employ in SEQC.
Bandic et al. (bandic2023mapping) map circuit partitioning to min-cut and use the QUBO model with quantum solvers to generate a set of cuts intended to minimize swaps between partitions. QUBO employs quantum optimization algorithms on quantum annealers to accelerate the solution process and remain practical, as the search space for transmission qubits is $2^{n}$. Instead, SEQC takes a “best-effort” approach that works well in practice, and does not require acceleration through quantum means.
Liu et al. (liu_2023_tackling) partition a circuit into blocks that span an identical number of qubits, perform permutation-aware synthesis on each block in parallel (solving for all possible qubit topologies), and finally use an augmented SABRE algorithm to map the blocks to the target device and route between blocks. SEQC takes a top-down approach instead, where SWAP permutations between modules are settled prior to any compilation tasks performed for each module (block). As a result, SEQC only needs to solve for one qubit layout per block, rather than all permutations.
Hungarian Qubit Assignment (HQA) (escofet2023hungarian) extends the Hungarian algorithm with the ability to adjust a qubit’s cost matrix according to behavior multiple time slices ahead. Pastor et al. (pastor2024circuit) use deep reinforcement learning to partition quantum circuits for execution across multiple modules. Both works aim to reduce inter-module swaps, but assume that all inter-module communications have the same cost, i.e., modules are fully connected, which is both impractical for large-scale systems and misses optimization opportunities. SEQC, in contrast, can work with arbitrary module topologies.
Ovide et al. (ovide2023mapping) leverage both quantum and classical methods for inter-module communication. While it advocates for a two-level approach to quantum circuit mapping, it leaves the actual problem of qubit placement and routing unsolved, which SEQC addresses with its algorithm.
8. Conclusion
[…]
Acknowledgements
This research was partially supported by NSF awards CNS-1763743 and CCF-2119069, Academic Year Undergraduate Research Award #2023-17704 by Northwestern University, and undergraduate summer research awards by the Robert R. McCormick School of Engineering and Applied Science and the Computer Science Department at Northwestern University. The authors thank Kabir Dubey for helping with earlier versions of this manuscript.
This research was supported in part through the computational resources and staff contributions provided for the Quest high performance computing facility at Northwestern University which is jointly supported by the Office of the Provost, the Office for Research, and Northwestern University Information Technology.
We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team.